### v2

# Improving Human-AI Coordination through Adversarial Training and Generative Models

Abstract
Being able to cooperate with new people is an important component of many economically valuable AI tasks, from household robotics to autonomous driving. However, generalizing to novel humans requires training on data that captures the diversity of human behaviors. Adversarial training is one avenue for searching for such data and ensuring that agents are robust. However, it is difficult to apply in the cooperative setting because adversarial policies intentionally learn to sabotage the task instead of simulating valid cooperation partners. To address this challenge, we propose a novel strategy for overcoming self-sabotage that combines a pre-trained generative model to simulate valid cooperative agent policies with adversarial training to maximize regret. We call our method GOAT: Generative Online Adversarial Training. In this framework, the GOAT dynamically searches for and generates coordination strategies where the learning policy—the Cooperator agent—underperforms. GOAT enables better generalization by exposing the Cooperator to various challenging interaction scenarios. We maintain realistic coordination strategies by updating only the generative model’s embedding while keeping its parameters frozen, thus avoiding adversarial exploitation. We evaluate GOAT with real human partners, and the results demonstrate state-of-the-art performance on the Overcooked benchmark, highlighting its effectiveness in generalizing to diverse human behaviors.
1 Introduction
Figure 1: Adversarial training framework for cooperative agents. There are two main components: (Left) A generative model encodes simulated agents into a latent space to learn diverse agent strategies, which are then used to simulate different types of agents. (Right) GOAT samples new agents to maximize the Cooperator agent’s regret, defined as the performance gap between self-play (GOAT sampled agents playing with themselves) and cross-play (the same GOAT sampled agent playing with the Cooperator). The key idea is that because the adversarial objective does not directly update the weights of the generative model, it can only simulate valid, cooperative partners that do not engage in self-sabotage. But by applying adversarial training to search over the policies that can be generated by the model, we can expose the Cooperator to a curriculum of challenging training partners, ensuring it is robust to playing with diverse human partners at test time.
In multi-agent human-AI cooperation (Carroll et al., 2020), training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge. AI systems must adapt to diverse, unpredictable, and often suboptimal human partners. Distribution shifts, limited training data, and complex human decision-making make training these systems difficult. Without adequate coverage of all possible human interactions, AI agents learn brittle policies and struggle to generalize to novel, unseen behaviors in real human cooperation partners.
Conventional approaches like behavior cloning (Carroll et al., 2020) and self-play (Samuel, 1959) are limited by training data, distribution shift issues, and a lack of coverage of the space of possible partner strategies. Population-Based Training (Jaderberg et al., 2017; Strouse et al., 2022; Zhao et al., 2022) has gained popularity as a zero-shot coordination technique, where a single Cooperator agent is trained against a population of simulated partners. To ensure this population covers the space of real human behaviors, many works have focused on enhancing population diversity (e.g. Hong et al. (2018); Lupu et al. (2021)). However, these techniques are often computationally expensive and still result in a poor match to the true distribution of human partners (Liang et al., 2024).
Adversarial training could provide a computationally efficient solution for training robust Cooperators by exposing both the Cooperator agent’s limitations and underexplored areas of the partner strategy space. Despite its potential, adversarial training within cooperative AI introduces unique challenges. For single-agent tasks, the Adversary’s sole objective is to challenge the learning agent as much as possible. However, in a collaborative task, the adversarial agent has to find weaknesses in the learner while still simulating a realistic cooperative partner. Suppose the Adversary is successful in finding weaknesses. In that case, it can become overly combative, resorting to strategies that sabotage the cooperation task, and is thus no longer a valid simulation of what a human cooperation partner might do. The challenge lies in simulating realistic yet challenging partners that maintain cooperative intentions while simultaneously identifying weaknesses in the AI agent’s policy.
For single-agent tasks, techniques like PAIRED (Dennis et al., 2021) constrain the Adversary so that it can only generate valid tasks for which there is a solution. This is accomplished by maximizing the learning agent’s regret, which is the performance gap between the optimal agent and the learning agent. The regret objective proves effective as the learning agent is actively challenged with a curriculum of increasingly difficult tasks, however, each task remains feasible.
In the cooperative setting, attempts have been made to create a regret-like objective for training more diverse simulated partner populations. Charakorn et al. (2023) proposes a training objective based on balancing the difference between two terms, self-play (SP) and cross-play (XP) performance. SP is the score a policy achieves when paired with itself, and XP is the score a policy receives when cooperating with a different policy. Adding the SP term ensures each policy can cooperate effectively (by achieving high scores with itself), while minimizing XP encourages the emergence of distinct strategies between policies. Thus, Charakorn et al. (2023) attempts to address the diversity problem using incompatibility with other policies as a proxy for diversity. However, naively minimizing XP can once again lead to policies learning to sabotage the game. They can learn a handshaking protocol, such that if their partner fails to complete the handshake correctly, they become adversarial and try to exploit weaknesses in the partner in order to minimize the XP score.
To overcome this issue, Sarkar et al. (2023) proposed balancing SP and XP optimization by adding "mixed-play" during the initial training phase, where initial game states are randomly sampled from both self-play and cross-play trajectories. Although mixed-play tries to address self-sabotage, it does not address the problem that minimizing cross-play inherently incentivizes destructive behavior. Further, mixed-play risks overfitting to previously learned conventions and limits how well the regret objective can be optimized.
How can we ensure that the adversarial agent can only generate valid cooperative behaviors that do not engage in self-sabotage but can still fully explore the space of challenging partner strategies? We propose a novel adversarial training method to achieve these criteria. Generative models like variational autoencoder (Kingma & Welling, 2022) enable novel data generation. Recent work from Liang et al. (2024) leverages this property of generative models to train a VAE on data generated from PBT methods like FCP (Strouse et al., 2022), CoMeDi (Sarkar et al., 2023), and MEP (Zhao et al., 2022) to be able to generate random diverse partners for training the Cooperator agent by sampling from the latent space. We leverage this method for simulating diverse partners in GOAT. We propose to place such a generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret (as approximated via SP-XP). Because we only update the weights of the Adversary and not the generative model, we ensure that all generated partners are valid and cooperative, thereby addressing the intentional sabotage seen in adversarial training. In this algorithm, the Adversary agent is motivated to maximize the regret, searching for valid coordination strategies where the Cooperator agent underperforms. In contrast, the Cooperator attempts to adapt to the Adversary and minimize the regret. Unlike previous adversarial approaches that risk learning degenerate strategies, this cyclic training method provides a natural curriculum for the Cooperator agent to train robustly and generalize to diverse partner strategies.
We evaluate GOAT using Overcooked (Carroll et al., 2020), a multi-agent cooperative cooking game where agents must collaborate to prepare soup recipes successfully. We tested GOAT live against real, novel human partners and found it significantly improved cooperation performance. GOAT achieved improved performance when evaluated against 5 competitive baselines (Zhao et al., 2022; Sarkar et al., 2023; Liang et al., 2024; Strouse et al., 2022; Carroll et al., 2020). The contributions of our paper—introducing an effective adversarial training method for human-AI cooperation and achieving state-of-the-art performance when evaluated in real-time with novel people—are described in detail on the cover page.
2 Related Work
Zero-Shot coordination (ZSC). Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings (Hu et al., 2020; Kirk et al., 2023). Training effective Cooperator agents requires navigating a vast coordination behavior space. Behavior cloning (Carroll et al., 2020) is limited by expensive and often scarce human data. Self-Play (SP) (Samuel, 1959; Silver et al., 2017; 2018; OpenAI et al., 2019) is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners. Population-Based Training (PBT) (Jaderberg et al., 2017; Vinyals et al., 2019; Parker-Holder et al., 2020; Jung et al., 2020) is another effective automatic training paradigm that has shown remarkable success in popular competitive games. PBT is widely adopted in the ZSC line of research (Hong et al., 2018; Strouse et al., 2022; Lupu et al., 2021; Charakorn et al., 2023; Zhao et al., 2022; Sarkar et al., 2023)
Population diversity objectives. In an effort to ensure the simulated agent population covers diverse human strategies, various methods (Hong et al., 2018; Parker-Holder et al., 2020; Wu et al., 2023; Lupu et al., 2021; Zhao et al., 2022; Charakorn et al., 2023; Sarkar et al., 2023; Rahman et al., 2023) improve PBT by optimizing the diversity of the agents in the population. The population can be created by reward-shaping (Tang et al., 2021; Yu et al., 2023), manual design (Ghosh et al., 2020; Wang et al., 2022; Xie et al., 2021), and quality diversity (Canaan et al., 2019; Pugh et al., 2016; Wu et al., 2023; Fontaine & Nikolaidis, 2021). However, they use domain knowledge and are not general and scalable. To automate the diverse population generation, two lines of work were introduced. Statistical methods (Hong et al., 2018; Eysenbach et al., 2018; Parker-Holder et al., 2020; Derek & Isola, 2021; Lupu et al., 2021; Zhao et al., 2022), where the entropy of action distributions or stationary trajectory distributions are maximized. However, Sarkar et al. (2023) demonstrated that trajectory distribution methods can result in minor variations in policies rather than behaviorally different policies. Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population (Cui et al., 2023; Charakorn et al., 2023; Sarkar et al., 2023). However, agents might learn to sabotage the cross-play game while retaining high self-play performance with handshaking protocols. To prevent overly adversarial behavior, Sarkar et al. (2023) introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both. However, such a solution still leads to rigid policies that do not generalize. In this work, we use online adversarial training to adapt to and exploit the cooperators’ weakness, rather than being limited by the coordination strategies in a pre-built population. And since we only sample from a frozen, pre-trained generative model, We prevent the adversary from learning handshaking protocols.
Adversarial training. Adversarial training (Tucker et al., 2020; Cui et al., 2023; Fujimoto & Pedersen, 2022) has emerged as one of the most effective methods for improving generalization to novel scenarios. Methods like PAIRED (Dennis et al., 2021) generate an automatic curriculum where an Adversary creates challenging yet feasible scenarios for the learning agent. This method is effective because it is based on a regret objective that incentivizes the Adversary agent to produce increasingly challenging but feasible tasks, addressing a crucial point: constraining the Adversary from over-exploiting the other agent’s weaknesses. GOAT builds upon these prior works by addressing the possible adversarial exploitation. We propose using regret-based adversarial training that naturally encourages the emergence of diverse strategic behaviors by forcing agents to exploit weaknesses in the Cooperator agent through actual competitive game outcomes.
Generative models for Cooperation. Derek & Isola (2021); Wang et al. (2023); Liang et al. (2024) show that generative models can be trained on policies to simulate diverse and valid collaborative policies. Recently, Liang et al. (2024) proposed training a Variational Autoencoder (Kingma & Welling, 2022), on simulated cooperative agent trajectories to generate synthetic training data. This approach outperforms previous methods because the generative model enables sampling novel partner strategies that interpolate between or combine strategies from the data. GOAT leverages this approach in an adversarial training loop where we adversarially sample from the latent space of the generative model to train the Cooperator agent robustly.
3 Preliminaries
Two-player Markov game: We consider the two-player Markov game (Littman, 1994) as a tuple $\bigl{(}\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr{)}$ where $\mathcal{S}$ is the state space shared by all the agents. $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ denote the action space of agents such that at each time step in the environment, each agent selects one action $a_{1,t}\in\mathcal{A}_{1}$, $a_{2,t}\in\mathcal{A}_{2}$ sampled from their own action space according to their policies $\pi_{1}:\mathcal{S}\to\Delta(\mathcal{A}_{1})$ and $\pi_{2}:\mathcal{S}\to\Delta(\mathcal{A}_{2})$. After taking these actions, they receive a shared immediate reward $r:\mathcal{S}\times\mathcal{A}_{1}\times\mathcal{A}_{2}\to\mathbb{R}$. The system then transitions to the next state $s_{t+1}$ according to the transition dynamics $\mathcal{T}(s_{t+1}\mid s_{t},a_{1,t},a_{2,t})\,\in\,\Delta(\mathcal{S})$ where $\Delta(\mathcal{S})$ denotes the set of probability distributions over $\mathcal{S}$. The parameter $\gamma\in[0,1)$ is the discount factor that weights immediate versus future rewards. Given two policies $\pi^{1}$ and $\pi^{2}$, mapping each state to a distribution over possible actions for Agents 1 and 2, respectively, the expected return is captured by the value function $V(\pi^{1},\pi^{2})\;=\;\mathbb{E}\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr{]}$ which sums the discounted rewards from each timestep. This formulation enables a systematic approach toward a cooperative decision-making algorithm between the two agents in a shared environment.
Generative model of cooperation strategies: Following Liang et al. (2024), we train a Variational Auto-Encoder (VAE) (Kingma & Welling, 2022) on diverse cooperative strategies. The VAE is trained to reconstruct trajectories from a dataset $D$ gathered using a PBT method. Here, we use Maximum Entropy Population (Zhao et al., 2022) since it was shown to provide the best performance in Liang et al. (2024). The VAE consists of an encoder $q(z|\tau;\phi)$, which maps a trajectory $\tau=(s_{0},a_{0},...,s_{T})$ to a latent space, where $\phi$ represents the parameters of the neural networks. The decoder $p(a_{t}|z,\tau_{0..t-1};\phi)$ predicts actions $a_{t}$ autoregressively using the latent $z$ and the agent’s history $\tau_{0..t-1}$. The VAE model is optimized using the ELBO loss described in (Kingma & Welling, 2022) as
$\mathcal{L}(\theta,\phi)=\mathbb{E}_{\tau\sim D}\Bigg{[}\mathbb{E}_{z\sim q(\cdot|\tau,\phi)}\left[\sum_{t=1}^{T}\log p(a_{t}|z,\tau_{0..t-1};\theta)\right]-\beta\cdot KL\big{(}q(z|\tau,\phi)\,\|\,\mathcal{N}(0,I)\big{)}\Bigg{]}$
The $\beta$ parameter balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space follows the prior that is a standard normal distribution $\mathcal{N}(0,I)$ for sampling new cooperation partners. The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations (Burgess et al., 2018). Thus, during training, the $\beta$ term was scheduled to automatically save different versions of KL-regularized VAE models. This method by Liang et al. (2024) has proved efficient in capturing diverse strategies and enabling strategy generation that is valid and cooperative.
4 GOAT: Generative Online Adversarial Training
We introduce GOAT, which combines generative modeling with online adversarial training to optimize regret-based objectives and improve human-AI coordination.
Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods (Charakorn et al., 2023). Here, an Adversary policy $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it. This type of optimization is widely known as minimax optimization in game theory.
$J_{XP}=\max_{\pi_{C}}\min_{\pi_{A}}V(\pi_{C},\pi_{A})$ (1)
As may be obvious, attempting to train agents to cooperate via XP minimization leads to a fundamentally misaligned optimization objective because it creates a conflict between agents to achieve the collective goal of maximizing common payoff. The Adversary agent is incentivized to develop behaviors that intentionally sabotage interactions rather than generate meaningful cooperation trajectories.
Generating adversarial cooperative agents. To address the above issues, we propose adding a pre-trained generative model (VAE) with frozen weights into the adversarial training loop. We hypothesize that adding a generative model that is trained on simulated agent coordination trajectories will ensure that adversarial training generates valid cooperative partners with no self-sabotage or handshaking issues. Previous XP minimization methods (Charakorn et al., 2023; Sarkar et al., 2023) represented policies as distinct, qualitative behaviors, requiring the algorithm to explore a combinatorial space of policies, whereas the VAE encodes the cooperation strategies as a continuous latent space allowing a smooth optimization landscape. We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation. Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce $z^{\prime}=\pi_{A}(z)$. This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$. Thus, the optimization problem is as follows:
$\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}\left[\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},\pi_{C}(a^{C}_{t}|\tau_{0:t-1}),p(a^{P}_{t}|\pi_{A}(z),\tau_{0:t-1};\theta))\right]$ (2)
$=$ $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}[V(\pi_{C},\pi_{P})]$ (3)
Note that this objective resolves the issue of directly training the partner policy that is choosing actions to minimize XP performance; Equation 3 reveals that the adversary agent does not have direct control over low-level actions, which instead are chosen by the frozen, pre-trained generative model. Instead, the adversary simply searches over the embedding space of the generative model to find challenging partners.
However, this minimax objective still does not guarantee effective exploration in the partner space as it optimizes for worst-case returns. Because the generative model is capable of simulating low-scoring, poor-quality partners, since these exist in the training data, the minimax objective will encourage the Adversary to focus on sampling these agents to minimize XP score. Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies, which the Cooperator policy is not able to perform well with.
Regret-based adversarial cooperative training. To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent to learn diverse strategies. We now formalize the terms for cooperative settings that we will use to formulate Regret.
Self-Play (SP): This play style, first used by Samuel (1959) to play checkers, is defined as the expected return when an agent is paired with itself:
$J_{SP}=\mathbb{E}[V(\pi_{P},\pi_{P})]$ (4)
Regret ($R$): As proposed initially by Robbins (1952), regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight. We formalize regret in the cooperative setting as the difference between the optimal self-play SP performance a policy can achieve with its training partner and the cross-play XP performance it achieves with another agent.
$Regret(\pi_{P},\pi_{C})=\mathbb{E}[V(\pi_{P},\pi_{P})-V(\pi_{P},\pi_{C})]=J_{SP}-J_{XP}$ (5)
Algorithm 1 GOAT
  Training VAE on trajectory dataset $D={\tau_{i}}_{i=1}^{N}$ where $\tau={s_{i},a_{i},r_{i}}_{i=1}^{N}$
  Initialize: VAE Decoder $p(a_{t}|z^{\prime},\tau_{0..t-1};\theta)$, Adversary $\pi_{A}$, Cooperator $\pi_{C}$
  while not converged do
     Sample $z\sim\mathcal{N}(0,I)$
     Map $z^{\prime}=\pi_{A}(z)$
     Generate $\pi_{P}=p(a_{t}|z^{\prime},\tau_{0..t-1};\theta)$
     Collect Partner self-play returns $J(\pi_{P},\pi_{P})$
     Collect Partner-Cooperator returns $J(\pi_{P},\pi_{C})$
     Compute $\text{REGRET}=J(\pi_{P},\pi_{P})-J(\pi_{P},\pi_{C})$
     Train Adversary Policy ($\pi_{A}$) with RL to maximize REGRET
     Train Cooperator Policy ($\pi_{C}$) with RL to minimize REGRET
  end while
At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$. This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$. The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns. Regret is evaluated after collecting the returns from both the game play styles. The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance. The full objective of the GOAT game is thus:
$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
The online, adversarial nature of this game ensures the Adversary is incentivized to continually challenge the Cooperator agent by discovering new partners it cannot play with yet, while the Cooperator is in turn corrects those deficiencies. As it does so, the regret decreases, and the Adversary is encouraged to explore the latent space to generate more challenging partners to maximize the regret. This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones. The GOAT is naturally incentivized to generate meaningful challenges. For invalid tasks, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no training signal to the Adversary. This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
5 Experiments
In this section, we will describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
Figure 2: Carroll et al. (2020) first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit. To increase coordination complexity, Liang et al. (2024) introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup. This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
Evaluation environment. We evaluate GOAT using the popular coordination benchmark Overcooked (Carroll et al., 2020), which requires two players to collaborate by dividing cooking tasks and avoiding obstructing each other. Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively. We test GOAT on the first five layouts introduced in Carroll et al. (2020), and add a more complex layout introduced by Liang et al. (2024), the Multi-Strategy Counter (all shown in Figure 2). Agents must follow small, intricate steps to prepare a dish successfully.
GOAT implementation. The Cooperator is trained using PPO, whereas the Adversary is method-independent, so we use REINFORCE to train it. Further details on the KL-regularization of the VAE and other architecture and hyperparameter details are available in the Appendix.
Baselines. We compare to prior baselines on human-AI coordination that are closely related to GOAT or are the previous state-of-the-art (SOTA). In particular, we evaluate GOAT against the conventional Behavior Cloning (BC) based method (Carroll et al., 2020) (BC+RL), population-based methods Fictitious Co-Play (FCP) (Strouse et al., 2022) and maximum entropy population (MEP) (Zhao et al., 2022), and GAMMA trained on MEP (MEP+GAMMA) Liang et al. (2024), which previously showed SOTA performance in real human evaluations.
Human Evaluation. We conducted a user study with 40 participants to discover which method best coordinated with humans and is evaluated as qualitatively better. Because human studies are expensive, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter. The latter presents a more complex cooperation task because partners can prepare two dishes using different ingredients and thus represents a more rigorous test of whether algorithms can generalize to cooperate with real humans. Participants were recruited from the Prolific online crowdsourcing platform following an IRB-approved protocol. Each participant was tasked with completing 6 total rounds of the Overcooked game with an AI agent partner, where each round presented a different agent model among one of the baselines or our method. In order to minimize trends due to gameplay order, the agent models are loaded in randomized order, with a randomly sampled checkpoint chosen from 1 of 5 random seeds. Additionally, each participant is instructed to answer qualitative questions on a 5-point Likert scale in order to rate their experience playing with each agent partner.
6 Experimental Results
Below, we described the results of empirical experiments organized around our central research questions.
RQ1: Will adversarial training result in a more robust Cooperator with better sample complexity? PBT methods, which rely on simulated populations of self-play partners, are often computationally expensive and have poor coverage of actual human behaviors. Training a Cooperator against them may not achieve good performance with people, and it may take a long time since the training does not focus on correcting weaknesses in the Cooperator but instead randomly samples from the population. We hypothesize that our adversarial training technique can more efficiently search the space of partners to make an agent that is more robust and requires less training time.
To validate this hypothesis, we investigate the learning curves of all simulated training methods, including our own, in Figure 3. Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work (Carroll et al., 2020; Strouse et al., 2022; Liang et al., 2024). The human proxy enables the automatic evaluation of agents’ performance throughout training.
Figure 3: Evaluation of GOAT against a simulated human policy trained with behavior cloning (BC). We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method (Liang et al., 2024). Error bars show the standard error over 5 random seeds. The Adversary is able to achieve better performance in fewer samples (RQ1).
As is evident in the curves, our adversarial training method achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques. We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations. Adversarial exploration of this space should enable rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies. By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to real human partners. Because it focuses training on weak points in the Cooperator’s policy, it creates an agent that is more robust to human stratgies, with lower sample complexity than traditional PBT methods.
Hypothesis Counter circuit ($p$-value) Multi-Strategy Counter ($p$-value)
CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
MEP $1.0$ $3.48\times 10^{-28}$
MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
[…]
RQ2: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans? As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-strategy Counter layouts. We recruited novel human players whom the agents had not previously seen. Figure 4 depicts the evaluation results for the layout and all the agents against human players. Each entry corresponds to a human-agent team. On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA see Table 1. We hypothesize this is because both methods are reaching close to optimal performance on this layout. However, on the more complex layout, Multi-Strategy Counter, the GOAT leads to significantly higher cooperation performance with real humans, leading to a 38% improvement over the previous state of the art.
Figure 4: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit and multi-strategy counter.
RQ3: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective? In Section 4, we described how the purely adversarial minimax objective (Equation 1) is misaligned with effectively training agents to cooperate. In Equation 3, we showed how integrating a generative model could help alleviate this problem. However, while the generative model can prevent sabotage behavior, that does not mean all agents it simulates will show productive cooperation behavior. We claimed that the purely adversarial minimax objective would still focus too much on poorly performing or incompetent partners, motivating the use of the regret objective in Equation 5. We now empirically support this claim by comparing GOAT to an ablation of our method, which uses the minimax objective in Equation 3, on the Multi-Strategy Counter layout. As we can see in Figure 6, the GOAT significantly outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT.
We further investigate the embeddings chosen by both types of Adversary in Figure 6 throughout the training period. We find that the minimax Adversary is pushed to a region (Red cluster in the plot) where the policies result in low reward behaviors. One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance but makes no attempts to cooperate. In the blue, we have embeddings chosen by the GOAT throughout the training period. We can observe that the regret objective enables the GOAT to explore the latent space effectively, thus covering a wider range of interesting behaviors.
Figure 5: Evaluation curves for comparison of performance of minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner. Error bars are the Standard Error of the Mean.
Figure 6: Adversary in red and GOAT in blue exploring the latent space over the training period.
RQ4: Will adversarial training create a curriculum of increasingly challenging strategies?
Figure 7: GOAT exploring the latent space exploration compared against GAMMA (Liang et al., 2024) in red over the whole training period.
The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges. As GOAT identifies underperforming coordination strategies, it should generate increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors. This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods. Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 7. We compare this to the distribution of embeddings the original GAMMA method (Liang et al., 2024) samples throughout training (red ball). We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple). As training progresses, it chooses a particular region to exploit. After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret. Finally, at the end of the training step in yellow, we can see that the GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
7 Conclusion
Our work demonstrates that combining generative modeling with adversarial training provides an effective approach for training AI agents that can coordinate with diverse human partners. By using a generative model to constrain the Adversary to valid cooperative behaviors while employing regret-based optimization to create increasingly challenging scenarios, we enable robust learning without the risk of degenerate strategies. Experimental results on the Overcooked benchmark show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners. The improved coverage of the partner space demonstrated by our agents suggests this approach successfully bridges the gap between training efficiency and the generalization to human behavior. Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training. From a broader perspective, this work represents a step toward AI systems that can flexibly coordinate with humans in real-world settings.
Broader Impact Statement
Our work on human-AI coordination aims to provide humanity with better cooperative AI that can adapt, understand, and generalize to diverse humans. While this could enhance collaborative robots, autonomous vehicles, and assistive technologies, our method specifically improves generalization to new partners without requiring extensive human data. GOAT has the potential to improve capabilities for household robots, vehicles, and collaborative AI systems where seamless human interaction adds value to people’s lives. This type of technology enhances human modelling, which can be deployed widely to help people from all walks of life and of all age groups. We would also like to note that the development of such an AI system will accelerate technological progress and requires responsible development and deployment.
8 Acknowledgements
This work was generously supported by the Amazon+UW Science Hub, the Cooperative AI Foundation, and the Sony Research Award Program.
Supplementary Materials
The following content was not necessarily subject to peer review.
9 Reproducibility
We have our demo, code, and extended experimental data on this webpage. https://sites.google.com/view/human-ai-goat-2025. We provide all the details to setup the environment and reproduce the results.
10 Implementation details
Generative models. The VAE model was trained on a dataset containing joint trajectories of two players. This dataset was created by evenly sampling from the pairs in the simulated agent population $\{\pi_{1},...,\pi_{N}\}$ generating 100k joint trajectories. The dataset was split in a 70/30 format for training and evaluation. The VAE was trained using ELBO loss, and linear scheduling of the KL penalty coefficient $\beta$ was applied to control the KL divergence of the posterior distribution. Automatically, checkpoints of VAE models of different strengths were saved.
Cooperator Agent The Cooperator agent was trained using PPO (Schulman et al., 2017). To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up. We adapted our codebase on the previous implementation of HSP (Yu et al., 2022) and GAMMA (Liang et al., 2024).
GOAT For the GOAT policy training, we usethe REINFORCE objective with mean and standard deviation bounds to keep the policy constrained in the VAE latent space. This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
11 Hyperparameters
The GOAT uses a simple policy and was trained using REINFORCE to generate latent vectors. The architecture and hyperparameters are defined below.
hyperparameter value
network architecture 4-layer MLP
hidden dimension 128
activation function ReLU
output dimension 2×2$\times$2 × z_dim
optimizer Adam
learning rate 0.001
weight decay 0.0001
latent dimension (z_dim) 16
KL coefficient 0.1
entropy coefficient 0.01
std scaling 4.0, 1.0
mean scaling 60.0
gradient clipping 1.0
learning rate scheduler start factor 1.0
learning rate scheduler end factor 0.9
Table 2: Hyperparameters for GOAT Generator
The Cooperator agent is trained using MAPPO with all overcooked layouts having similar low-level implementation details like architecture and hyperparameters. Every policy network has the same structure, with a CNN coming after an RNN (we use GRU).
hyperparameter value
CNN kernels [3, 3], [3, 3], [3, 3]
CNN channels [32, 64, 32]
hidden layer size [64]
recurrent layer size 64
activation function ReLU
weight decay 0
environment steps 100M (simulated data) or 150M (human data)
parallel environments 200
episode length 400
PPO batch size $2\times 200\times 400$
PPO epoch 15
PPO learning rate 0.0005
Generalized Advantage Estimator (GAE) $\lambda$ 0.95
discounting factor $\gamma$ 0.99
Table 3: Policy hyperparameters
The architecture of the policy model and the generative model are comparable. The representations are transformed into variational posterior and action reconstruction predictions using an encoder head and a decoder head.
hyperparameter value
CNN kernels [3, 3], [3, 3], [3, 3]
CNN channels [32, 64, 32]
hidden layer size [256]
recurrent layer size 256
activation function ReLU
weight decay 0.0001
parallel environments 200
episode length 400
epoch 500
chunk length 100
learning rate 0.0005
KL penalty coefficient $\beta$ $0\to 1$
latent variable dimension 16
Table 4: Hyperparameters for VAE models
12 Computational Resources
Our primary tests were carried out on AMD EPYC 64-Core Processor and NVIDIA L40s/L40 clusters. One Cooperator agent can be trained in roughly a day. The primary tests require roughly 3600 GPU hours. We perform some preliminary experiments to determine the best training frameworks and hyperparameters.

### v3

@@
-Improving Human-AI Coordination through Adversarial Training and Generative Models
+Improving Human-AI Coordination through Online Adversarial Training and Generative Models
 Abstract
@@
 However, generalizing to novel humans requires training on data that captures the diversity of human behaviors.
-Adversarial training is one avenue for searching for such data and ensuring that agents are robust.
-However, it is difficult to apply in the cooperative setting because adversarial policies intentionally learn to sabotage the task instead of simulating valid cooperation partners.
-To address this challenge, we propose a novel strategy for overcoming self-sabotage that combines a pre-trained generative model to simulate valid cooperative agent policies with adversarial training to maximize regret.
+Adversarial training is a promising method that allows dynamic data generation and ensures that agents are robust.
+It creates a feedback loop where the agent’s performance influences the generation of new adversarial data, which can be used immediately to train the agent.
+However, adversarial training is difficult to apply in a cooperative task; how can we train an adversarial cooperator?
+We propose a novel strategy that combines a pre-trained generative model to simulate valid cooperative agent policies with adversarial training to maximize regret.
 We call our method GOAT: Generative Online Adversarial Training.
-In this framework, the GOAT dynamically searches for and generates coordination strategies where the learning policy—the Cooperator agent—underperforms.
+In this framework, the GOAT dynamically searches the latent space of the generative model for coordination strategies where the learning policy—the Cooperator agent—underperforms.
 GOAT enables better generalization by exposing the Cooperator to various challenging interaction scenarios.
-We maintain realistic coordination strategies by updating only the generative model’s embedding while keeping its parameters frozen, thus avoiding adversarial exploitation.
+We maintain realistic coordination strategies by keeping the generative model frozen, thus avoiding adversarial exploitation.
 We evaluate GOAT with real human partners, and the results demonstrate state-of-the-art performance on the Overcooked benchmark, highlighting its effectiveness in generalizing to diverse human behaviors.
+[^1] Please check out our live interactive demo, game play videos, and training code at our website.
+We also provide a runnable Google Colab for the cooperative matrix game.
 1 Introduction
@@
 But by applying adversarial training to search over the policies that can be generated by the model, we can expose the Cooperator to a curriculum of challenging training partners, ensuring it is robust to playing with diverse human partners at test time.
-In multi-agent human-AI cooperation (Carroll et al., 2020), training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge.
-AI systems must adapt to diverse, unpredictable, and often suboptimal human partners.
-Distribution shifts, limited training data, and complex human decision-making make training these systems difficult.
+In multi-agent human-AI cooperation [3], training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge.
+Distribution shift, limited training data, and complex human decision-making make training these systems difficult.
 Without adequate coverage of all possible human interactions, AI agents learn brittle policies and struggle to generalize to novel, unseen behaviors in real human cooperation partners.
-Conventional approaches like behavior cloning (Carroll et al., 2020) and self-play (Samuel, 1959) are limited by training data, distribution shift issues, and a lack of coverage of the space of possible partner strategies.
-Population-Based Training (Jaderberg et al., 2017; Strouse et al., 2022; Zhao et al., 2022) has gained popularity as a zero-shot coordination technique, where a single Cooperator agent is trained against a population of simulated partners.
-To ensure this population covers the space of real human behaviors, many works have focused on enhancing population diversity (e.g.
-Hong et al.
-(2018); Lupu et al.
-(2021)).
-However, these techniques are often computationally expensive and still result in a poor match to the true distribution of human partners (Liang et al., 2024).
 Adversarial training could provide a computationally efficient solution for training robust Cooperators by exposing both the Cooperator agent’s limitations and underexplored areas of the partner strategy space.
-Despite its potential, adversarial training within cooperative AI introduces unique challenges.
-For single-agent tasks, the Adversary’s sole objective is to challenge the learning agent as much as possible.
+It can be used to automatically generate a curriculum of challenging tasks targeted to the learner’s weaknesses, resulting in a more robust policy [6].
+Yet despite its potential, adversarial training within cooperative AI introduces unique challenges; what does it mean to train an adversarial cooperator?
+In a single-agent task, the Adversary has the sole objective to challenge the learning agent as much as possible by minimizing its performance.
 However, in a collaborative task, the adversarial agent has to find weaknesses in the learner while still simulating a realistic cooperative partner.
@@
 In that case, it can become overly combative, resorting to strategies that sabotage the cooperation task, and is thus no longer a valid simulation of what a human cooperation partner might do.
-The challenge lies in simulating realistic yet challenging partners that maintain cooperative intentions while simultaneously identifying weaknesses in the AI agent’s policy.
-For single-agent tasks, techniques like PAIRED (Dennis et al., 2021) constrain the Adversary so that it can only generate valid tasks for which there is a solution.
+The challenge lies in simulating realistic yet challenging partners that maintain cooperative intentions while simultaneously correcting weaknesses in the AI agent’s policy.
+For single-agent tasks, techniques like PAIRED [6] constrain the Adversary so that it can only generate valid tasks for which there is a solution.
 This is accomplished by maximizing the learning agent’s regret, which is the performance gap between the optimal agent and the learning agent.
@@
 Charakorn et al.
-(2023) proposes a training objective based on balancing the difference between two terms, self-play (SP) and cross-play (XP) performance.
+[4] proposes a training objective to make a population of mutually incompatible partners, based on balancing the difference between two terms, self-play (SP) and cross-play (XP) performance.
 SP is the score a policy achieves when paired with itself, and XP is the score a policy receives when cooperating with a different policy.
 Adding the SP term ensures each policy can cooperate effectively (by achieving high scores with itself), while minimizing XP encourages the emergence of distinct strategies between policies.
-Thus, Charakorn et al.
-(2023) attempts to address the diversity problem using incompatibility with other policies as a proxy for diversity.
 However, naively minimizing XP can once again lead to policies learning to sabotage the game.
-They can learn a handshaking protocol, such that if their partner fails to complete the handshake correctly, they become adversarial and try to exploit weaknesses in the partner in order to minimize the XP score.
-To overcome this issue, Sarkar et al.
-(2023) proposed balancing SP and XP optimization by adding "mixed-play" during the initial training phase, where initial game states are randomly sampled from both self-play and cross-play trajectories.
-Although mixed-play tries to address self-sabotage, it does not address the problem that minimizing cross-play inherently incentivizes destructive behavior.
-Further, mixed-play risks overfitting to previously learned conventions and limits how well the regret objective can be optimized.
+They can learn a handshaking protocol, such that if their partner fails to complete the handshake correctly, they become adversarial and try to strategically sabotage the game to minimize the XP score.
+While attempts have been made to add additional objectives to counterbalance this effect Sarkar et al.
+[28], they have not been able to address the issue that minimizing cross-play inherently incentivizes uncooperative behavior.
 How can we ensure that the adversarial agent can only generate valid cooperative behaviors that do not engage in self-sabotage but can still fully explore the space of challenging partner strategies?
 We propose a novel adversarial training method to achieve these criteria.
-Generative models like variational autoencoder (Kingma & Welling, 2022) enable novel data generation.
-Recent work from Liang et al.
-(2024) leverages this property of generative models to train a VAE on data generated from PBT methods like FCP (Strouse et al., 2022), CoMeDi (Sarkar et al., 2023), and MEP (Zhao et al., 2022) to be able to generate random diverse partners for training the Cooperator agent by sampling from the latent space.
-We leverage this method for simulating diverse partners in GOAT.
-We propose to place such a generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret (as approximated via SP-XP).
+We first pre-train a generative model of partner strategies on valid, diverse cooperation data, following Liang et al.
+[18].
+We then place this generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret (as approximated via SP-XP).
 Because we only update the weights of the Adversary and not the generative model, we ensure that all generated partners are valid and cooperative, thereby addressing the intentional sabotage seen in adversarial training.
@@
 Unlike previous adversarial approaches that risk learning degenerate strategies, this cyclic training method provides a natural curriculum for the Cooperator agent to train robustly and generalize to diverse partner strategies.
-We evaluate GOAT using Overcooked (Carroll et al., 2020), a multi-agent cooperative cooking game where agents must collaborate to prepare soup recipes successfully.
-We tested GOAT live against real, novel human partners and found it significantly improved cooperation performance.
-GOAT achieved improved performance when evaluated against 5 competitive baselines (Zhao et al., 2022; Sarkar et al., 2023; Liang et al., 2024; Strouse et al., 2022; Carroll et al., 2020).
-The contributions of our paper—introducing an effective adversarial training method for human-AI cooperation and achieving state-of-the-art performance when evaluated in real-time with novel people—are described in detail on the cover page.
+We evaluate GOAT on a toy environment, One-Step Cooperative Matrix Game (CMG), and Overcooked [3], a popular human-AI zero shot coordination benchmark, which is a cooperative cooking game where agents must collaborate to prepare soup recipes successfully.
+On the CMG toy environment, we find that GOAT discovers and exploits all the regions of reward space with multiple maxima and minima, while other baselines [43, 20, 28] get stuck or improperly assign weights to certain modes.
+We tested GOAT live against real, novel human partners and found that it significantly improves cooperation performance.
+GOAT achieves improved performance when evaluated against 5 competitive baselines [43, 28, 18, 32, 3].
+The contributions of this paper are as follows:
+  • We introduce a novel adversarial training method for zero-shot coordination, which uses an Adversary to search over the space of partner policies that a generative model can simulate.
+Because the generative model can only simulate valid cooperative partners, it ensures the adversarial training process generates challenging but realistic cooperative partners.
+  • We train both the Adversary and Cooperator agents online in a common payoff game.
+This results in a dynamic curriculum where the Adversary actively adopts strategies tailored to challenge the Cooperator policy’s current weaknesses, while the Cooperator learns to effectively correct them.
+  • We perform a real-time evaluation with novel human partners, using the popular Overcooked benchmark [3].
+As compared to recent high-scoring techniques, we show that our method leads to the best cooperation performance with diverse people.
 2 Related Work
 Zero-Shot coordination (ZSC).
-Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings (Hu et al., 2020; Kirk et al., 2023).
+Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings [13, 17].
 Training effective Cooperator agents requires navigating a vast coordination behavior space.
-Behavior cloning (Carroll et al., 2020) is limited by expensive and often scarce human data.
-Self-Play (SP) (Samuel, 1959; Silver et al., 2017; 2018; OpenAI et al., 2019) is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners.
-Population-Based Training (PBT) (Jaderberg et al., 2017; Vinyals et al., 2019; Parker-Holder et al., 2020; Jung et al., 2020) is another effective automatic training paradigm that has shown remarkable success in popular competitive games.
-PBT is widely adopted in the ZSC line of research (Hong et al., 2018; Strouse et al., 2022; Lupu et al., 2021; Charakorn et al., 2023; Zhao et al., 2022; Sarkar et al., 2023)
+Behavior cloning [3] is limited by expensive and often scarce human data.
+Self-Play (SP) [27, 30, 31, 21] is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners.
+Population-Based Training (PBT) [14, 35, 22, 15] has become a popular automatic training paradigm for ZSC, in which a Cooperator policy is trained using a large population of simulated partners [12, 32, 20, 4, 43, 28].
 Population diversity objectives.
-In an effort to ensure the simulated agent population covers diverse human strategies, various methods (Hong et al., 2018; Parker-Holder et al., 2020; Wu et al., 2023; Lupu et al., 2021; Zhao et al., 2022; Charakorn et al., 2023; Sarkar et al., 2023; Rahman et al., 2023) improve PBT by optimizing the diversity of the agents in the population.
-The population can be created by reward-shaping (Tang et al., 2021; Yu et al., 2023), manual design (Ghosh et al., 2020; Wang et al., 2022; Xie et al., 2021), and quality diversity (Canaan et al., 2019; Pugh et al., 2016; Wu et al., 2023; Fontaine & Nikolaidis, 2021).
-However, they use domain knowledge and are not general and scalable.
-To automate the diverse population generation, two lines of work were introduced.
-Statistical methods (Hong et al., 2018; Eysenbach et al., 2018; Parker-Holder et al., 2020; Derek & Isola, 2021; Lupu et al., 2021; Zhao et al., 2022), where the entropy of action distributions or stationary trajectory distributions are maximized.
-However, Sarkar et al.
-(2023) demonstrated that trajectory distribution methods can result in minor variations in policies rather than behaviorally different policies.
-Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population (Cui et al., 2023; Charakorn et al., 2023; Sarkar et al., 2023).
+In an effort to ensure the simulated agent population covers diverse human strategies, various methods [12, 22, 39, 20, 43, 4, 28, 24] improve PBT by optimizing the diversity of the agents in the population.
+The population can be created by reward-shaping [33, 42], manual design [11, 37, 40], and quality diversity [2, 23, 39, 9].
+However, these methods typically use domain knowledge and are not general and scalable.
+Statistical methods [12, 8, 22, 7, 20, 43] automate the diverse population generation but Sarkar et al.
+[28] demonstrated that trajectory distribution methods can result in minor variations in the policies rather than behaviorally different policies.
+Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population [5, 4, 28].
 However, agents might learn to sabotage the cross-play game while retaining high self-play performance with handshaking protocols.
 To prevent overly adversarial behavior, Sarkar et al.
-(2023) introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both.
+[28] introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both.
 However, such a solution still leads to rigid policies that do not generalize.
 In this work, we use online adversarial training to adapt to and exploit the cooperators’ weakness, rather than being limited by the coordination strategies in a pre-built population.
-And since we only sample from a frozen, pre-trained generative model, We prevent the adversary from learning handshaking protocols.
+And since we only sample from a frozen, pre-trained generative model, we prevent the adversary from learning handshaking protocols or sabotaging behavior.
+Generative models for Cooperation.
+Derek and Isola [7], Wang et al.
+[36], Liang et al.
+[18] show that generative models can be trained on policies to simulate diverse and valid collaborative policies.
+Recently, Liang et al.
+[18] proposed training a Variational Autoencoder [16], on simulated cooperative agent trajectories to generate synthetic training data to improve human-AI coordination.
+This approach outperforms previous methods because the generative model enables sampling novel partner strategies that interpolate between or combine strategies from the data.
+GOAT leverages this approach in an adversarial training loop.
 Adversarial training.
-Adversarial training (Tucker et al., 2020; Cui et al., 2023; Fujimoto & Pedersen, 2022) has emerged as one of the most effective methods for improving generalization to novel scenarios.
-Methods like PAIRED (Dennis et al., 2021) generate an automatic curriculum where an Adversary creates challenging yet feasible scenarios for the learning agent.
-This method is effective because it is based on a regret objective that incentivizes the Adversary agent to produce increasingly challenging but feasible tasks, addressing a crucial point: constraining the Adversary from over-exploiting the other agent’s weaknesses.
-GOAT builds upon these prior works by addressing the possible adversarial exploitation.
-We propose using regret-based adversarial training that naturally encourages the emergence of diverse strategic behaviors by forcing agents to exploit weaknesses in the Cooperator agent through actual competitive game outcomes.
-Generative models for Cooperation.
-Derek & Isola (2021); Wang et al.
-(2023); Liang et al.
-(2024) show that generative models can be trained on policies to simulate diverse and valid collaborative policies.
-Recently, Liang et al.
-(2024) proposed training a Variational Autoencoder (Kingma & Welling, 2022), on simulated cooperative agent trajectories to generate synthetic training data.
-This approach outperforms previous methods because the generative model enables sampling novel partner strategies that interpolate between or combine strategies from the data.
-GOAT leverages this approach in an adversarial training loop where we adversarially sample from the latent space of the generative model to train the Cooperator agent robustly.
+Adversarial training [34, 5, 10, 26] has emerged as one of the most effective methods for improving generalization to novel scenarios.
+Methods like PAIRED [6] generate an automatic curriculum where an Adversary creates challenging yet feasible scenarios for the learning agent.
+This method is effective because it is based on a regret objective that incentivizes the Adversary agent to produce increasingly challenging but feasible tasks, addressing a crucial point: constraining the Adversary from minimizing performance to the point where the learner cannot receive any reward and training becomes stalled.
+We propose using regret-based adversarial training to focus on presenting partner strategies that target weaknesses in the Cooperator agent, but are still valid cooperative strategies.
+Our results demonstrate that the use of a regret-based objective is critical to the success of our method.
 3 Preliminaries
-Two-player Markov game: We consider the two-player Markov game (Littman, 1994) as a tuple $\bigl{(}\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr{)}$ where $\mathcal{S}$ is the state space shared by all the agents.
+Two-player Markov game: We consider the two-player Markov game [19] as a tuple $\bigl{(}\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr{)}$ where $\mathcal{S}$ is the state space shared by all the agents.
 $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ denote the action space of agents such that at each time step in the environment, each agent selects one action $a_{1,t}\in\mathcal{A}_{1}$, $a_{2,t}\in\mathcal{A}_{2}$ sampled from their own action space according to their policies $\pi_{1}:\mathcal{S}\to\Delta(\mathcal{A}_{1})$ and $\pi_{2}:\mathcal{S}\to\Delta(\mathcal{A}_{2})$.
@@
 The parameter $\gamma\in[0,1)$ is the discount factor that weights immediate versus future rewards.
-Given two policies $\pi^{1}$ and $\pi^{2}$, mapping each state to a distribution over possible actions for Agents 1 and 2, respectively, the expected return is captured by the value function $V(\pi^{1},\pi^{2})\;=\;\mathbb{E}\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr{]}$ which sums the discounted rewards from each timestep.
-This formulation enables a systematic approach toward a cooperative decision-making algorithm between the two agents in a shared environment.
+Given two policies $\pi_{1}$ and $\pi_{2}$, mapping each state to a distribution over possible actions for Agents 1 and 2, respectively, the expected return is captured by the value function $V(\pi_{1},\pi_{2})\;=\;\mathbb{E}\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr{]}$ which sums the discounted rewards from each timestep.
+We will use $V(\pi_{1},\pi_{2})$ in later sections to discuss how agents can maximize their joint return.
 Generative model of cooperation strategies: Following Liang et al.
-(2024), we train a Variational Auto-Encoder (VAE) (Kingma & Welling, 2022) on diverse cooperative strategies.
-The VAE is trained to reconstruct trajectories from a dataset $D$ gathered using a PBT method.
-Here, we use Maximum Entropy Population (Zhao et al., 2022) since it was shown to provide the best performance in Liang et al.
-(2024).
-The VAE consists of an encoder $q(z|\tau;\phi)$, which maps a trajectory $\tau=(s_{0},a_{0},...,s_{T})$ to a latent space, where $\phi$ represents the parameters of the neural networks.
+[18], we train a Variational Auto-Encoder (VAE) [16] on diverse cooperative strategies.
+The VAE is trained to reconstruct coordination trajectories from a dataset $D$, which is obtained by using simulated PBT agents to interact with each other in the environment.
+Here, we use Maximum Entropy Population [43] to provide the simulated partner agents, since MEP was shown to provide the best performance in Liang et al.
+[18].
+After producing the dataset $D$ using MEP agents, we then train the generative model, specifically the VAE, on this data.
+The VAE consists of an encoder $q(z|\tau;\phi)$, which maps a trajectory $\tau=(s_{0},a_{0},...,s_{T})$ to a latent space, where $\phi$ represents the parameters of the decoder.
 The decoder $p(a_{t}|z,\tau_{0..t-1};\phi)$ predicts actions $a_{t}$ autoregressively using the latent $z$ and the agent’s history $\tau_{0..t-1}$.
-The VAE model is optimized using the ELBO loss described in (Kingma & Welling, 2022) as
-$\mathcal{L}(\theta,\phi)=\mathbb{E}_{\tau\sim D}\Bigg{[}\mathbb{E}_{z\sim q(\cdot|\tau,\phi)}\left[\sum_{t=1}^{T}\log p(a_{t}|z,\tau_{0..t-1};\theta)\right]-\beta\cdot KL\big{(}q(z|\tau,\phi)\,\|\,\mathcal{N}(0,I)\big{)}\Bigg{]}$
-The $\beta$ parameter balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space follows the prior that is a standard normal distribution $\mathcal{N}(0,I)$ for sampling new cooperation partners.
-The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations (Burgess et al., 2018).
-Thus, during training, the $\beta$ term was scheduled to automatically save different versions of KL-regularized VAE models.
+The VAE model is optimized using the ELBO loss described in [16] as
+$\mathcal{L}(\theta,\phi)=\mathbb{E}_{\tau\sim D}\Bigg{[}\mathbb{E}_{z\sim q(\cdot|\tau,\phi)}\left[\sum_{t=1}^{T}\log p(a_{t}|z,\tau_{0..t-1};\phi)\right]-\beta\cdot KL\big{(}q(z|\tau,\phi)\,\|\,\mathcal{N}(0,I)\big{)}\Bigg{]}$
+The $\beta$ parameter balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space resembles a standard normal distribution $\mathcal{N}(0,I)$.
+This is important, because we can then sample new cooperation partners by sampling a $z\sim\mathcal{N}(0,I)$, and using the decoder to obtain a partner policy $\pi_{P}:=p(a_{t}|z,\tau_{0..t-1};\phi)$.
+The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations [1].
+During training, the $\beta$ term was scheduled to automatically save different versions of KL-regularized VAE models.
 This method by Liang et al.
-(2024) has proved efficient in capturing diverse strategies and enabling strategy generation that is valid and cooperative.
+[18] has proved efficient in capturing diverse strategies and enabling strategy generation that is valid and cooperative.
 4 GOAT: Generative Online Adversarial Training
-We introduce GOAT, which combines generative modeling with online adversarial training to optimize regret-based objectives and improve human-AI coordination.
-Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods (Charakorn et al., 2023).
+We introduce GOAT, which combines generative modeling with online regret-based adversarial training to enable robust human-AI coordination.
+Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods [4].
 Here, an Adversary policy $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it.
-This type of optimization is widely known as minimax optimization in game theory.
+This type of optimization is widely known as minimax optimization in game theory:
 $J_{XP}=\max_{\pi_{C}}\min_{\pi_{A}}V(\pi_{C},\pi_{A})$ (1)
-As may be obvious, attempting to train agents to cooperate via XP minimization leads to a fundamentally misaligned optimization objective because it creates a conflict between agents to achieve the collective goal of maximizing common payoff.
+As may be obvious, attempting to train agents to cooperate via XP minimization leads to a fundamentally misaligned optimization objective because it conflicts with the cooperative objective of maximizing common payoff.
 The Adversary agent is incentivized to develop behaviors that intentionally sabotage interactions rather than generate meaningful cooperation trajectories.
@@
 To address the above issues, we propose adding a pre-trained generative model (VAE) with frozen weights into the adversarial training loop.
-We hypothesize that adding a generative model that is trained on simulated agent coordination trajectories will ensure that adversarial training generates valid cooperative partners with no self-sabotage or handshaking issues.
-Previous XP minimization methods (Charakorn et al., 2023; Sarkar et al., 2023) represented policies as distinct, qualitative behaviors, requiring the algorithm to explore a combinatorial space of policies, whereas the VAE encodes the cooperation strategies as a continuous latent space allowing a smooth optimization landscape.
-We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation.
-Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce $z^{\prime}=\pi_{A}(z)$.
+We hypothesize that adding a generative model trained solely on cooperative policy data will ensure that adversarial training generates only valid cooperative partners.
+By design, the generative model is not capable of simulating self-sabotage or handshaking behaviors.
+Instead, we train a separate Adversary policy to search for embedding vectors $z$ that decode into partners that minimize XP performance for the specific Cooperator agent we seek to train.
+The Adversary can thus search over a well-structured, continuous latent space, allowing a smooth optimization landscape.
+We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation 1.
+Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce an Adversary embedding $z^{\prime}=\pi_{A}(z)$.
 This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$.
@@
 $=$ $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}[V(\pi_{C},\pi_{P})]$ (3)
-Note that this objective resolves the issue of directly training the partner policy that is choosing actions to minimize XP performance; Equation 3 reveals that the adversary agent does not have direct control over low-level actions, which instead are chosen by the frozen, pre-trained generative model.
-Instead, the adversary simply searches over the embedding space of the generative model to find challenging partners.
 However, this minimax objective still does not guarantee effective exploration in the partner space as it optimizes for worst-case returns.
-Because the generative model is capable of simulating low-scoring, poor-quality partners, since these exist in the training data, the minimax objective will encourage the Adversary to focus on sampling these agents to minimize XP score.
-Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies, which the Cooperator policy is not able to perform well with.
+Therefore, the Adversary will only search for the worst agents that the generative model can simulate.
+Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies which could have good performance, but the Cooperator failed to cooperate well with them.
 Regret-based adversarial cooperative training.
-To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent to learn diverse strategies.
-We now formalize the terms for cooperative settings that we will use to formulate Regret.
-Self-Play (SP): This play style, first used by Samuel (1959) to play checkers, is defined as the expected return when an agent is paired with itself:
+To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent.
+We now formalize the definition of Regret we will use for cooperative settings.
+Self-Play (SP): This play style, first used by Samuel [27] to play checkers, is defined as the expected return when an agent is paired with itself:
 $J_{SP}=\mathbb{E}[V(\pi_{P},\pi_{P})]$ (4)
-Regret ($R$): As proposed initially by Robbins (1952), regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight.
-We formalize regret in the cooperative setting as the difference between the optimal self-play SP performance a policy can achieve with its training partner and the cross-play XP performance it achieves with another agent.
+Regret ($R$): As proposed initially by Robbins [25], regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight.
+We formalize regret in the cooperative setting as the performance gap between the cross-play (XP) performance an agent achieves with a partner $\pi_{P}$, and $\pi_{P}$’s optimal self-play (SP) performance.
 $Regret(\pi_{P},\pi_{C})=\mathbb{E}[V(\pi_{P},\pi_{P})-V(\pi_{P},\pi_{C})]=J_{SP}-J_{XP}$ (5)
+At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$.
+This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$.
+The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns.
+Regret is evaluated after collecting the returns from both games.
+The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance.
+The full objective of the GOAT game is thus:
+$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
+The online optimization as shown in algorithm 1 ensures that the Adversary is incentivized to continually discover new partner strategies that are challenging for the Cooperator, by searching the latent space to maximize regret.
+The Cooperator in turn learns to adapt to these partner strategies by minimizing the regret.
+This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones.
+For invalid partners, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no incentive to the Adversary.
+This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
 Algorithm 1 GOAT
-  Training VAE on trajectory dataset $D={\tau_{i}}_{i=1}^{N}$ where $\tau={s_{i},a_{i},r_{i}}_{i=1}^{N}$
+  Training VAE on trajectory dataset $D={\tau_{i}}_{i=1}^{N}$ where $\tau=\{s_{i},a_{i},r_{i}\}_{i=1}^{N}$
   Initialize: VAE Decoder $p(a_{t}|z^{\prime},\tau_{0..t-1};\theta)$, Adversary $\pi_{A}$, Cooperator $\pi_{C}$
@@
   end while
-At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$.
-This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$.
-The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns.
-Regret is evaluated after collecting the returns from both the game play styles.
-The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance.
-The full objective of the GOAT game is thus:
-$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
-The online, adversarial nature of this game ensures the Adversary is incentivized to continually challenge the Cooperator agent by discovering new partners it cannot play with yet, while the Cooperator is in turn corrects those deficiencies.
-As it does so, the regret decreases, and the Adversary is encouraged to explore the latent space to generate more challenging partners to maximize the regret.
-This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones.
-The GOAT is naturally incentivized to generate meaningful challenges.
-For invalid tasks, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no training signal to the Adversary.
-This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
 5 Experiments
-In this section, we will describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
+In this section, we describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
+We evaluate GOAT using two popular coordination benchmarks: 1)1)1 ) One-Step Cooperative Matrix Game (CMG) [20, 43, 4] and 2)2)2 ) Overcooked [3].
 Figure 2: Carroll et al.
-(2020) first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit.
+[3] first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit.
 To increase coordination complexity, Liang et al.
-(2024) introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup.
+[18] introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup.
 This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
-Evaluation environment.
-We evaluate GOAT using the popular coordination benchmark Overcooked (Carroll et al., 2020), which requires two players to collaborate by dividing cooking tasks and avoiding obstructing each other.
+5.1 One-Step Cooperative Matrix Game (CMG):
+This is a single-step game where two agents (Row agent and Column agent) independently select an action by choosing a row and a column, respectively, and the intersection of their choices returns the associated reward according to the payoff matrix.
+Both agents receive the same reward and attempt to maximize the common payoff.
+A typical payoff matrix in CMG is defined by $(M,\{k_{m}\},\{r_{m}\})$ where $M$ is number of solutions, $k_{m}$ is number of compatible actions, and $r_{m}$ is the reward associated with each solution.
+We use a challenging multimodal layout called CMG-M as shown in Figure 3.
+5.2 Overcooked:
+This is a collaborative game where two players need to collaborate by dividing cooking tasks and avoiding obstructing each other.
 Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively.
-We test GOAT on the first five layouts introduced in Carroll et al.
-(2020), and add a more complex layout introduced by Liang et al.
-(2024), the Multi-Strategy Counter (all shown in Figure 2).
-Agents must follow small, intricate steps to prepare a dish successfully.
-GOAT implementation.
-The Cooperator is trained using PPO, whereas the Adversary is method-independent, so we use REINFORCE to train it.
-Further details on the KL-regularization of the VAE and other architecture and hyperparameter details are available in the Appendix.
-Baselines.
-We compare to prior baselines on human-AI coordination that are closely related to GOAT or are the previous state-of-the-art (SOTA).
-In particular, we evaluate GOAT against the conventional Behavior Cloning (BC) based method (Carroll et al., 2020) (BC+RL), population-based methods Fictitious Co-Play (FCP) (Strouse et al., 2022) and maximum entropy population (MEP) (Zhao et al., 2022), and GAMMA trained on MEP (MEP+GAMMA) Liang et al.
-(2024), which previously showed SOTA performance in real human evaluations.
-Human Evaluation.
+We test GOAT on the first five layouts introduced in [3], and add a more complex layout introduced by Liang et al.
+[18], the Multi-Strategy Counter (all shown in Figure 2).
+5.3 GOAT implementation & Baselines:
+The Cooperator is trained using PPO, and the Adversary is trained using REINFORCE.
+Training the Adversary is RL-algorithm-agnostic as it does not depend on state and is a one-step optimization process.
+Further details on the KL-regularization of the VAE, training procedure for GOAT, and other architecture and hyperparameters are available in the Appendix.
+We compare to 5 competitive prior baselines on human-AI coordination: Behavior Cloning (BC) partners with RL cooperator [3] (BC+RL), PBT methods Fictitious Co-Play (FCP) [32], maximum entropy population (MEP) [43], CoMeDi [28], and GAMMA trained on MEP (MEP+GAMMA) [18], which previously showed state-of-the-art performance in real human evaluations.
+5.4 Human Evaluation:
 We conducted a user study with 40 participants to discover which method best coordinated with humans and is evaluated as qualitatively better.
-Because human studies are expensive, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter.
+Because we want the most challenging tasks with the most diverse strategy space, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter.
 The latter presents a more complex cooperation task because partners can prepare two dishes using different ingredients and thus represents a more rigorous test of whether algorithms can generalize to cooperate with real humans.
@@
 In order to minimize trends due to gameplay order, the agent models are loaded in randomized order, with a randomly sampled checkpoint chosen from 1 of 5 random seeds.
-Additionally, each participant is instructed to answer qualitative questions on a 5-point Likert scale in order to rate their experience playing with each agent partner.
 6 Experimental Results
-Below, we described the results of empirical experiments organized around our central research questions.
-RQ1: Will adversarial training result in a more robust Cooperator with better sample complexity?
+(a) CMG Layout
+(b) Self-play
+(c) TrajeDi
+(d) MEP
+(e) CoMeDi
+(f) GOAT
+Figure 3: To test the ability of the method to discover and exploit multiple minima and maxima in the problem space, we introduce a custom CMG layout (a).
+This reward structure is designed to have three scenarios: no reward, suboptimal rewards, and global maximum reward.
+The policies for the methods were aggregated and normalized into a heatmap.
+6.1 RQ1: Can GOAT discover and exploit multiple cooperative modes?
+For GOAT to be effective, the VAE should be able to discover and exploit regions of the reward space.
+To test this, we use a one-step cooperative matrix game (CMG).
+We test the cooperative policy generation capability of VAE by training it on trajectory data generated by the population of CoMeDi agents.
+Since CoMeDi data is cooperative, we hypothesize that the VAE should only generate policies that are in distribution.
+To keep the evaluation fair, we used a population size of 8 for all the methods tested on CMG-M.
+During training, GOAT covers all the modes and correctly assigns importance to the agent’s actions.
+As seen in Figure 3, the baselines self-play, Trajedi, and MEP get stuck and might not discover other modes.
+CoMeDi discovers all the regions and gives the highest priority to the global maximum, but spreads probability mass roughly equally across all other modes, regardless of their payoff.
+In contrast, GOAT not only covers all modes but is better at assigning higher probability to more rewarding modes (for example, assigning low probability to modes 4 and 7, which have low payoff).
+We note that the quality of VAE training and the amount of training data greatly influence the resulting GOAT policy.
+To robustly train the VAE, we follow Liang et al.
+[18] and augment the 8 policy pool data by pairing the row and column agents of one policy with other policies.
+In our experiments, we saw that higher $\beta$ for KL regularization in VAE resulted in better generation of actions.
+6.2 RQ2: Will adversarial training result in a more robust Cooperator with better sample complexity?
 PBT methods, which rely on simulated populations of self-play partners, are often computationally expensive and have poor coverage of actual human behaviors.
@@
 To validate this hypothesis, we investigate the learning curves of all simulated training methods, including our own, in Figure 4.
-Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work (Carroll et al., 2020; Strouse et al., 2022; Liang et al., 2024).
+Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work [3, 32, 18].
 The human proxy enables the automatic evaluation of agents’ performance throughout training.
+As is evident in the curves, GOAT achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques.
+We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations.
+Adversarial exploration of this space enables rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies.
+By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to diverse partners.
+Because it efficiently targets training to weak points in the Cooperator’s policy, it has lower sample complexity than traditional PBT methods.
 Figure 4: Evaluation of GOAT against a simulated human policy trained with behavior cloning (BC).
-We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method (Liang et al., 2024).
+We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method, GAMMA [18].
 Error bars show the standard error over 5 random seeds.
-The Adversary is able to achieve better performance in fewer samples (RQ1).
-As is evident in the curves, our adversarial training method achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques.
-We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations.
-Adversarial exploration of this space should enable rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies.
-By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to real human partners.
-Because it focuses training on weak points in the Cooperator’s policy, it creates an agent that is more robust to human stratgies, with lower sample complexity than traditional PBT methods.
-Hypothesis Counter circuit ($p$-value) Multi-Strategy Counter ($p$-value)
-CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
-FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
-BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
-MEP $1.0$ $3.48\times 10^{-28}$
-MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
-[…]
-RQ2: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans?
+6.3 RQ3: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans?
 As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-strategy Counter layouts.
 We recruited novel human players whom the agents had not previously seen.
-Figure 4 depicts the evaluation results for the layout and all the agents against human players.
-Each entry corresponds to a human-agent team.
-On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA see Table 1.
+Figure 5 depicts the evaluation results for the two layouts and all the agents against human players.
+On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA (see Table 4 for statistical significance results).
 We hypothesize this is because both methods are reaching close to optimal performance on this layout.
 However, on the more complex layout, Multi-Strategy Counter, the GOAT leads to significantly higher cooperation performance with real humans, leading to a 38% improvement over the previous state of the art.
-Figure 4: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit and multi-strategy counter.
-RQ3: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective?
+To test these differences for yourself, please visit our webpage (https://sites.google.com/view/goat-2025/), which provides an interactive demo in which you can play Overcooked with both GOAT and the baselines.
+Figure 5: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit (left plot) and multi-strategy counter (right plot).
+6.4 RQ4: Will adversarial training create a curriculum of increasingly challenging strategies by discovering them in latent space?
+The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges.
+GOAT identifies underperforming coordination strategies and generates increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors.
+This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods.
+Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
+Figure 6: Left: GOAT exploring the latent space exploration compared against GAMMA [18] in red over the whole training period.
+Right: Minimax Adversary in red and GOAT in blue exploring the latent space over the training period.
+We provide a visual example related to this on our webpage.
+The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 6.
+We compare this to the distribution of embeddings of the original GAMMA method [18] samples throughout training (red ball).
+We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple).
+As training progresses, it chooses a particular region to exploit.
+After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret.
+Finally, at the end of the training step in yellow, GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
+6.5 RQ5: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective?
 In Section 4, we described how the purely adversarial minimax objective (Equation 1) is misaligned with effectively training agents to cooperate.
@@
 We now empirically support this claim by comparing GOAT to an ablation of our method, which uses the minimax objective in Equation 3, on the Multi-Strategy Counter layout.
-As we can see in Figure 6, the GOAT significantly outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT.
-We further investigate the embeddings chosen by both types of Adversary in Figure 6 throughout the training period.
-We find that the minimax Adversary is pushed to a region (Red cluster in the plot) where the policies result in low reward behaviors.
-One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance but makes no attempts to cooperate.
+As we can see in Figure 7, the GOAT outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT.
+We further investigate the embeddings chosen by both types of Adversary throughout the training period in Figure 6 (right).
+We find that the minimax Adversary mostly fixates on a single region (red cluster in the plot) where the policies result in low reward behaviors.
+One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance, but makes no attempts to cooperate.
 In the blue, we have embeddings chosen by the GOAT throughout the training period.
-We can observe that the regret objective enables the GOAT to explore the latent space effectively, thus covering a wider range of interesting behaviors.
-Figure 5: Evaluation curves for comparison of performance of minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner.
-Error bars are the Standard Error of the Mean.
-Figure 6: Adversary in red and GOAT in blue exploring the latent space over the training period.
-RQ4: Will adversarial training create a curriculum of increasingly challenging strategies?
-Figure 7: GOAT exploring the latent space exploration compared against GAMMA (Liang et al., 2024) in red over the whole training period.
-The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges.
-As GOAT identifies underperforming coordination strategies, it should generate increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors.
-This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods.
-Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
-The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 7.
-We compare this to the distribution of embeddings the original GAMMA method (Liang et al., 2024) samples throughout training (red ball).
-We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple).
-As training progresses, it chooses a particular region to exploit.
-After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret.
-Finally, at the end of the training step in yellow, we can see that the GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
+We can observe that the regret objective causes the GOAT to move to explore multiple modes over the course of training, exploring the latent space more effectively and thus covering a wider range of interesting behaviors.
 7 Conclusion
 Our work demonstrates that combining generative modeling with adversarial training provides an effective approach for training AI agents that can coordinate with diverse human partners.
-By using a generative model to constrain the Adversary to valid cooperative behaviors while employing regret-based optimization to create increasingly challenging scenarios, we enable robust learning without the risk of degenerate strategies.
+Using a generative model and regret-based optimization to constrain the Adversary enables GOAT to generate valid, meaningful cooperation partners while posing a curriculum of increasingly challenging scenarios, adapted to the weaknesses of the Cooperator’s policy.
+Thus, we enable robust learning without the risk of degenerate strategies.
 Experimental results on the Overcooked benchmark show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners.
@@
 Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training.
-From a broader perspective, this work represents a step toward AI systems that can flexibly coordinate with humans in real-world settings.
-Broader Impact Statement
-Our work on human-AI coordination aims to provide humanity with better cooperative AI that can adapt, understand, and generalize to diverse humans.
-While this could enhance collaborative robots, autonomous vehicles, and assistive technologies, our method specifically improves generalization to new partners without requiring extensive human data.
-GOAT has the potential to improve capabilities for household robots, vehicles, and collaborative AI systems where seamless human interaction adds value to people’s lives.
-This type of technology enhances human modelling, which can be deployed widely to help people from all walks of life and of all age groups.
-We would also like to note that the development of such an AI system will accelerate technological progress and requires responsible development and deployment.
 8 Acknowledgements
-This work was generously supported by the Amazon+UW Science Hub, the Cooperative AI Foundation, and the Sony Research Award Program.
-Supplementary Materials
-The following content was not necessarily subject to peer review.
-9 Reproducibility
-We have our demo, code, and extended experimental data on this webpage. https://sites.google.com/view/human-ai-goat-2025. We provide all the details to setup the environment and reproduce the results.
-10 Implementation details
+This work was generously supported by the Cooperative AI Foundation, UW-Amazon Science Gift Hub, Amazon Middle Mile Products and Technology Team, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), and Microsoft Accelerate Foundation Models Research Program.
+Appendix A Reproducibility
+We have our live interactive demo and extended experimental data on this webpage.
+https://sites.google.com/view/goat-2025.
+We provide all the details to set up the environment and reproduce the results.
+For the One-Step Cooperative Matrix Game please find Google Colab with code and all the instructions to run the code.
+Appendix B Implementation details
 Generative models.
 The VAE model was trained on a dataset containing joint trajectories of two players.
-This dataset was created by evenly sampling from the pairs in the simulated agent population $\{\pi_{1},...,\pi_{N}\}$ generating 100k joint trajectories.
+This dataset was created by evenly sampling from the pairs in the simulated agent population $\{\pi_{1},...,\pi_{N}\}$, generating 100k joint trajectories.
 The dataset was split in a 70/30 format for training and evaluation.
@@
 Automatically, checkpoints of VAE models of different strengths were saved.
-Cooperator Agent The Cooperator agent was trained using PPO (Schulman et al., 2017).
+For CMG, the training data used for VAE was 8 policies generated by CoMeDi on CMG.
+Due to a smaller amount of data, we used data augmentation to train the VAE.
+Cooperator Agent The Cooperator agent was trained using PPO [29].
 To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up.
-We adapted our codebase on the previous implementation of HSP (Yu et al., 2022) and GAMMA (Liang et al., 2024).
-GOAT For the GOAT policy training, we usethe REINFORCE objective with mean and standard deviation bounds to keep the policy constrained in the VAE latent space.
+We adapted our codebase on the previous implementation of HSP [41] and GAMMA [18].
+GOAT For the GOAT policy training, we use the REINFORCE objective with KL divergence penalty for drifting away from the prior, mean, and standard deviation hard bounds to keep the policy constrained in the VAE latent space.
 This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
-11 Hyperparameters
+Appendix C Hyperparameters
 The GOAT uses a simple policy and was trained using REINFORCE to generate latent vectors. The architecture and hyperparameters are defined below.
@@
 The architecture of the policy model and the generative model are comparable.
-The representations are transformed into variational posterior and action reconstruction predictions using an encoder head and a decoder head.
+The representations are transformed into a variational posterior and action reconstruction predictions using an encoder head and a decoder head.
 hyperparameter value
@@
 Table 3: Hyperparameters for VAE models
-12 Computational Resources
+Appendix D Computational Resources
 Our primary tests were carried out on AMD EPYC 64-Core Processor and NVIDIA L40s/L40 clusters.
@@
 We perform some preliminary experiments to determine the best training frameworks and hyperparameters.
+Appendix E Minimax and Regret comparison on overcooked
+Figure 7: Performance comparison between minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner.
+Error bars are the Standard Error of the Mean.
+Appendix F Cooperative strategies learned by VAE
+Figure 8: Left: We collect 4 cooperative policies on CMG-M using CoMeDi and then aggregate them.
+Right: The policies are used for training the VAE, where the latent dimension is 2 and we estimate the policy reconstruction from the VAE.
+Figure 9: After training a VAE on a 4-population pool of CoMeDi agents, we visualize the generation by interpolating the dimensions of the latent vector.
+Particularly for this example, we choose the latent dimension to be 2.
+We can see that VAE trained on cooperative policies has only encoded those policies in the latent space.
+As we move from one side of the plot to the other, we can see that different cells are activated, and it covers the cooperative policies as seen in the Figure 8
+Appendix G Bonferroni Tests
+Hypothesis Counter circuit ($p$-value) Multi-Strategy Counter ($p$-value)
+CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
+FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
+BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
+MEP $1.0$ $3.48\times 10^{-28}$
+MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
+[…]
+Appendix H Tests on CMG-S
+(a) CMG-S layout
+(b) Self-play
+(c) TrajeDi
+(d) MEP
+(e) CoMeDi
+(f) GOAT
+Figure 10: We provide additional tests of the CMG-S problem space.
+In the problem space, rewards flow in a direction that creates a smooth learning landscape for the algorithms to exploit.
+The policies for the methods were aggregated and normalized into a heatmap.
+GOAT successfully assigns priorities to regions.
+Appendix I Broader Impact Statement
+Our work on human-AI coordination aims to provide humanity with better cooperative AI that can adapt, understand, and generalize to diverse humans.
+While this could enhance collaborative robots, autonomous vehicles, and assistive technologies, our method specifically improves generalization to new partners without requiring extensive human data.
+GOAT has the potential to improve capabilities for household robots, vehicles, and collaborative AI systems where seamless human interaction adds value to people’s lives.
+This type of technology enhances human modelling, which can be deployed widely to help people from all walks of life and of all age groups.
+We would also like to note that the development of such an AI system will accelerate technological progress and requires responsible development and deployment.
+Appendix J Human qualitative ratings.
+(a) Counter circuit layout
+(b) Multi-strategy Counter
+Figure 11: Human evaluations of agent behavior.
+Participants rated their agreement with the statement "The agent adapted to me when making decisions" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree.
+GOAT shows decent adaptability on simpler layout, whereas it appears more adaptable on harder layout.
+(a) Counter circuit layout
+(b) Multi-strategy Counter
+Figure 12: User assessment of agent coordination.
+Participants rated their agreement with the statement "The agent’s actions were human-like" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree.
+GOAT is seen as more confident in coordinating on harder layouts.
+(a) Counter circuit layout
+(b) Multi-strategy Counter
+Figure 13: User assessments of agent frustration levels.
+Participants rated their agreement with the statement "The agent’s behavior was frustrating" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree.
+GOAT receives comparable ratings for showing less frustrating behavior compared to their respective baselines.
+Appendix K Limitation
+The effectiveness of this method heavily depends on the training of VAE models and enough data available for VAE to learn the distribution instead of overfitting to a few examples.
+It would be a promising method for fields like robotics and human-AI interaction, where data availability is not an issue.
+We also found that in sparse reward problems, this method discovers solutions and trains the cooperator robustly.

### v4

@@
 Abstract
-Being able to cooperate with new people is an important component of many economically valuable AI tasks, from household robotics to autonomous driving.
+Being able to cooperate with diverse humans is an important component of many economically valuable AI tasks, from household robotics to autonomous driving.
 However, generalizing to novel humans requires training on data that captures the diversity of human behaviors.
@@
 We call our method GOAT: Generative Online Adversarial Training.
-In this framework, the GOAT dynamically searches the latent space of the generative model for coordination strategies where the learning policy—the Cooperator agent—underperforms.
+In this framework, GOAT dynamically searches the latent space of the generative model for coordination strategies where the learning policy—the Cooperator agent—underperforms.
 GOAT enables better generalization by exposing the Cooperator to various challenging interaction scenarios.
@@
 We evaluate GOAT with real human partners, and the results demonstrate state-of-the-art performance on the Overcooked benchmark, highlighting its effectiveness in generalizing to diverse human behaviors.
-[^1] Please check out our live interactive demo, game play videos, and training code at our website.
-We also provide a runnable Google Colab for the cooperative matrix game.
+[^1] Please check out our live interactive demo, game play videos, and training code at our website https://sites.google.com/view/goat-2025/home
 1 Introduction
+In multi-agent human-AI cooperation (Carroll et al., 2020), training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge.
+Distribution shift, limited training data, and highly dynamic human decision-making make training these systems difficult.
+Without adequate coverage of all possible human interactions, AI agents learn brittle policies and struggle to generalize to novel, unseen behaviors in real human cooperation partners.
 Figure 1: Adversarial training framework for cooperative agents.
-There are two main components: (Left) A generative model encodes simulated agents into a latent space to learn diverse agent strategies, which are then used to simulate different types of agents.
-(Right) GOAT samples new agents to maximize the Cooperator agent’s regret, defined as the performance gap between self-play (GOAT sampled agents playing with themselves) and cross-play (the same GOAT sampled agent playing with the Cooperator).
-The key idea is that because the adversarial objective does not directly update the weights of the generative model, it can only simulate valid, cooperative partners that do not engage in self-sabotage.
-But by applying adversarial training to search over the policies that can be generated by the model, we can expose the Cooperator to a curriculum of challenging training partners, ensuring it is robust to playing with diverse human partners at test time.
-In multi-agent human-AI cooperation [3], training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge.
-Distribution shift, limited training data, and complex human decision-making make training these systems difficult.
-Without adequate coverage of all possible human interactions, AI agents learn brittle policies and struggle to generalize to novel, unseen behaviors in real human cooperation partners.
-Adversarial training could provide a computationally efficient solution for training robust Cooperators by exposing both the Cooperator agent’s limitations and underexplored areas of the partner strategy space.
-It can be used to automatically generate a curriculum of challenging tasks targeted to the learner’s weaknesses, resulting in a more robust policy [6].
+(Left) A generative model encodes simulated agents into a latent space to learn diverse agent strategies, which are then used to generate different types of training partners.
+(Right) GOAT samples new partners to maximize the Cooperator agent’s regret, defined as the performance gap between self-play (the partner playing with itself) and cross-play (the partner playing with the Cooperator).
+The key idea is that the adversarial objective is constrained by the frozen generative model, which prevents it from generating self-sabotaging partners.
+But by applying regret-based adversarial training to search over the policies that can be generated by the model, we can expose the Cooperator to a curriculum of challenging training partners, ensuring it is robust to interacting with diverse human partners at test time.
+Adversarial training (Tucker et al., 2020; Cui et al., 2023; Fujimoto & Pedersen, 2022; Rutherford et al., 2024) could provide a computationally efficient solution for training robust Cooperators by exposing both the Cooperator agent’s limitations and underexplored areas of the partner strategy space.
+It can be used to automatically generate a curriculum of challenging tasks targeted to the learner’s weaknesses, resulting in a more robust policy Dennis et al.
+(2021).
 Yet despite its potential, adversarial training within cooperative AI introduces unique challenges; what does it mean to train an adversarial cooperator?
-In a single-agent task, the Adversary has the sole objective to challenge the learning agent as much as possible by minimizing its performance.
-However, in a collaborative task, the adversarial agent has to find weaknesses in the learner while still simulating a realistic cooperative partner.
-Suppose the Adversary is successful in finding weaknesses.
-In that case, it can become overly combative, resorting to strategies that sabotage the cooperation task, and is thus no longer a valid simulation of what a human cooperation partner might do.
+In a single-agent task, Adversary’s sole objective is to challenge the learning agent as much as possible by minimizing its performance.
+However, in a collaborative task, the adversarial agent has to find weaknesses in the learner while still simulating a realistic cooperative partner policy.
+If the adversary is not constrained to such policies, then it can become overly combative, resorting to strategies that sabotage the cooperation task, and is thus no longer a valid simulation of what a human cooperation partner might do.
 The challenge lies in simulating realistic yet challenging partners that maintain cooperative intentions while simultaneously correcting weaknesses in the AI agent’s policy.
-For single-agent tasks, techniques like PAIRED [6] constrain the Adversary so that it can only generate valid tasks for which there is a solution.
+For single-agent tasks, techniques like Unsupervised Environment Design (UED) (Dennis et al., 2021; Parker-Holder et al., 2022; Jiang et al., 2021) constrains the Adversary so that it can only generate valid tasks for which there is a solution.
 This is accomplished by maximizing the learning agent’s regret, which is the performance gap between the optimal agent and the learning agent.
-The regret objective proves effective as the learning agent is actively challenged with a curriculum of increasingly difficult tasks, however, each task remains feasible.
-In the cooperative setting, attempts have been made to create a regret-like objective for training more diverse simulated partner populations.
-Charakorn et al.
-[4] proposes a training objective to make a population of mutually incompatible partners, based on balancing the difference between two terms, self-play (SP) and cross-play (XP) performance.
-SP is the score a policy achieves when paired with itself, and XP is the score a policy receives when cooperating with a different policy.
-Adding the SP term ensures each policy can cooperate effectively (by achieving high scores with itself), while minimizing XP encourages the emergence of distinct strategies between policies.
-However, naively minimizing XP can once again lead to policies learning to sabotage the game.
-They can learn a handshaking protocol, such that if their partner fails to complete the handshake correctly, they become adversarial and try to strategically sabotage the game to minimize the XP score.
+The regret objective proves effective because it challenges the learning agent with a curriculum of increasingly difficult, yet still feasible tasks.
+In the cooperative setting, attempts have been made to create a regret-like objective (Charakorn et al., 2023; Villin et al., 2025; Rutherford et al., 2024; Wang et al., 2025; Erlebach & Cook, 2024).
+In methods like Charakorn et al.
+(2023), a diverse population is trained by optimizing for the difference between self-play and cross-play performance.
+However, naively applying this type of adversarial objective can lead to policies learning to sabotage the game.
 While attempts have been made to add additional objectives to counterbalance this effect Sarkar et al.
-[28], they have not been able to address the issue that minimizing cross-play inherently incentivizes uncooperative behavior.
-How can we ensure that the adversarial agent can only generate valid cooperative behaviors that do not engage in self-sabotage but can still fully explore the space of challenging partner strategies?
+(2023), they have not been able to address the issue that minimizing cross-play inherently incentivizes overly adversarial, uncooperative behavior.
+How can we ensure that the adversarial agent can only generate cooperative behaviors that do not engage in self-sabotage but can still fully explore the space of challenging partner strategies?
 We propose a novel adversarial training method to achieve these criteria.
-We first pre-train a generative model of partner strategies on valid, diverse cooperation data, following Liang et al.
-[18].
-We then place this generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret (as approximated via SP-XP).
-Because we only update the weights of the Adversary and not the generative model, we ensure that all generated partners are valid and cooperative, thereby addressing the intentional sabotage seen in adversarial training.
-In this algorithm, the Adversary agent is motivated to maximize the regret, searching for valid coordination strategies where the Cooperator agent underperforms.
+We first pre-train a generative model of partner strategies on diverse cooperation data, following Liang et al.
+(2024).
+We then place this generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret.
+Because we only update the weights of the Adversary and not the generative model, we ensure that all generated partners are cooperative, thereby addressing the intentional sabotage seen in adversarial training.
+In GOAT, the Adversary is motivated to maximize the regret, searching for valid coordination strategies where the Cooperator agent underperforms.
 In contrast, the Cooperator attempts to adapt to the Adversary and minimize the regret.
-Unlike previous adversarial approaches that risk learning degenerate strategies, this cyclic training method provides a natural curriculum for the Cooperator agent to train robustly and generalize to diverse partner strategies.
-We evaluate GOAT on a toy environment, One-Step Cooperative Matrix Game (CMG), and Overcooked [3], a popular human-AI zero shot coordination benchmark, which is a cooperative cooking game where agents must collaborate to prepare soup recipes successfully.
-On the CMG toy environment, we find that GOAT discovers and exploits all the regions of reward space with multiple maxima and minima, while other baselines [43, 20, 28] get stuck or improperly assign weights to certain modes.
-We tested GOAT live against real, novel human partners and found that it significantly improves cooperation performance.
-GOAT achieves improved performance when evaluated against 5 competitive baselines [43, 28, 18, 32, 3].
+Unlike previous adversarial approaches that risk learning degenerate strategies, this cyclic training method provides a natural curriculum for the Cooperator agent and unlocks the benefits of adversarial training for robustly generalizing to cooperate with diverse partners.
+We evaluate GOAT on three distinct domains: 1) One-Step Cooperative Matrix Game (CMG); 2) Cooperative Reaching Game (CRG); and 3) Overcooked (Carroll et al., 2020), a popular human-AI zero-shot coordination benchmark.
+We compare to 5 competitive baselines (Zhao et al., 2022; Sarkar et al., 2023; Liang et al., 2024; Strouse et al., 2022; Carroll et al., 2020) and show that GOAT achieves improved performance across all three popular environments.
+In the Overcooked environment, we tested GOAT live against real, novel human partners and found that it significantly improves cooperation performance.
 The contributions of this paper are as follows:
-  • We introduce a novel adversarial training method for zero-shot coordination, which uses an Adversary to search over the space of partner policies that a generative model can simulate.
-Because the generative model can only simulate valid cooperative partners, it ensures the adversarial training process generates challenging but realistic cooperative partners.
-  • We train both the Adversary and Cooperator agents online in a common payoff game.
-This results in a dynamic curriculum where the Adversary actively adopts strategies tailored to challenge the Cooperator policy’s current weaknesses, while the Cooperator learns to effectively correct them.
-  • We perform a real-time evaluation with novel human partners, using the popular Overcooked benchmark [3].
+  • We introduce a novel adversarial training method for zero-shot coordination that trains both the Adversary and Cooperator agents online in a common payoff game, resulting in a dynamic curriculum where the Adversary searches over the space of partner policies that a generative model can simulate.
+The generative model is limited to simulating cooperative partners, which ensures the adversarial training process generates challenging but realistic cooperative partners.
+  • We outperform 5 competitive baselines across 3 distinct popular cooperative environments.
+  • We perform a real-time evaluation with novel human partners, using the popular Overcooked benchmark (Carroll et al., 2020) commonly used in human-AI cooperation research like Strouse et al.
+(2022); Charakorn et al.
+(2023); Zhao et al.
+(2022); Sarkar et al.
+(2023); Liang et al.
+(2024).
 As compared to recent high-scoring techniques, we show that our method leads to the best cooperation performance with diverse people.
@@
 Zero-shot coordination (ZSC).
-Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings [13, 17].
-Training effective Cooperator agents requires navigating a vast coordination behavior space.
-Behavior cloning [3] is limited by expensive and often scarce human data.
-Self-Play (SP) [27, 30, 31, 21] is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners.
-Population-Based Training (PBT) [14, 35, 22, 15] has become a popular automatic training paradigm for ZSC, in which a Cooperator policy is trained using a large population of simulated partners [12, 32, 20, 4, 43, 28].
+Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings, and thus navigate a vast coordination behavior space.
+(Hu et al., 2020; Kirk et al., 2023).
+Behavior cloning (Carroll et al., 2020) is limited by expensive and often scarce human data.
+Self-Play (SP) (Samuel, 1959; Silver et al., 2017; 2018; OpenAI et al., 2019) is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners.
+Population-Based Training (PBT) (Jaderberg et al., 2017; Vinyals et al., 2019; Parker-Holder et al., 2020; Jung et al., 2020) has become a popular automatic training paradigm for ZSC, in which a Cooperator policy is trained using a large population of simulated partners (Hong et al., 2018; Strouse et al., 2022; Lupu et al., 2021; Charakorn et al., 2023; Zhao et al., 2022; Sarkar et al., 2023).
 Population diversity objectives.
-In an effort to ensure the simulated agent population covers diverse human strategies, various methods [12, 22, 39, 20, 43, 4, 28, 24] improve PBT by optimizing the diversity of the agents in the population.
-The population can be created by reward-shaping [33, 42], manual design [11, 37, 40], and quality diversity [2, 23, 39, 9].
-However, these methods typically use domain knowledge and are not general and scalable.
-Statistical methods [12, 8, 22, 7, 20, 43] automate the diverse population generation but Sarkar et al.
-[28] demonstrated that trajectory distribution methods can result in minor variations in the policies rather than behaviorally different policies.
-Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population [5, 4, 28].
-However, agents might learn to sabotage the cross-play game while retaining high self-play performance with handshaking protocols.
-To prevent overly adversarial behavior, Sarkar et al.
-[28] introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both.
-However, such a solution still leads to rigid policies that do not generalize.
-In this work, we use online adversarial training to adapt to and exploit the cooperators’ weakness, rather than being limited by the coordination strategies in a pre-built population.
-And since we only sample from a frozen, pre-trained generative model, we prevent the adversary from learning handshaking protocols or sabotaging behavior.
+In an effort to ensure the simulated agent population covers diverse human strategies, various methods (Hong et al., 2018; Parker-Holder et al., 2020; Wu et al., 2023; Lupu et al., 2021; Zhao et al., 2022; Charakorn et al., 2023; Sarkar et al., 2023; Rahman et al., 2023; Yuan et al., 2023) improve PBT by optimizing the diversity of the agents in the population.
+The population can be created by reward-shaping (Tang et al., 2021; Yu et al., 2023), manual design (Ghosh et al., 2020; Wang et al., 2022; Xie et al., 2021), and quality diversity (Canaan et al., 2019; Pugh et al., 2016; Wu et al., 2023; Fontaine & Nikolaidis, 2021).
+However, these methods typically use domain knowledge and fail to generalize and scale.
+Statistical methods like (Hong et al., 2018; Eysenbach et al., 2018; Parker-Holder et al., 2020; Derek & Isola, 2021; Lupu et al., 2021; Zhao et al., 2022) automate the diverse population generation, but Sarkar et al.
+(2023) demonstrated that trajectory distribution methods can result in minor variations in the policies rather than behaviorally different policies.
+Min-Max optimization.
+Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population (Cui et al., 2023; Charakorn et al., 2023; Rahman et al., 2023; 2024; Villin et al., 2025; Sarkar et al., 2023).
+However, agents might learn to sabotage the cross-play game while retaining high self-play performance with handshaking protocols, such that if their partner fails to complete the handshake correctly, they throw the game.
+To address this, Sarkar et al.
+(2023) introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both.
+However, such a solution still leads to rigid policies that fail to generalize, and does not address the fundamental problem that training a partner to minimize cross-play incentives sabotage.
+In this work, we only sample partners from a frozen, pre-trained generative model; since we do not train the partner policy on the adversarial objective, we prevent handshaking and sabotage.
+In addition, unlike prior work on this topic, we use online adversarial training to find and exploit weaknesses in the cooperator, instead of using adversarial objectives to pre-train a diverse partner population.
 Generative models for Cooperation.
-Derek and Isola [7], Wang et al.
-[36], Liang et al.
-[18] show that generative models can be trained on policies to simulate diverse and valid collaborative policies.
+Derek & Isola (2021); Wang et al.
+(2023); Liang et al.
+(2024) show that generative models can be trained on policies to simulate diverse and valid collaborative policies.
 Recently, Liang et al.
-[18] proposed training a Variational Autoencoder [16], on simulated cooperative agent trajectories to generate synthetic training data to improve human-AI coordination.
-This approach outperforms previous methods because the generative model enables sampling novel partner strategies that interpolate between or combine strategies from the data.
-GOAT leverages this approach in an adversarial training loop.
-Adversarial training.
-Adversarial training [34, 5, 10, 26] has emerged as one of the most effective methods for improving generalization to novel scenarios.
-Methods like PAIRED [6] generate an automatic curriculum where an Adversary creates challenging yet feasible scenarios for the learning agent.
-This method is effective because it is based on a regret objective that incentivizes the Adversary agent to produce increasingly challenging but feasible tasks, addressing a crucial point: constraining the Adversary from minimizing performance to the point where the learner cannot receive any reward and training becomes stalled.
-We propose using regret-based adversarial training to focus on presenting partner strategies that target weaknesses in the Cooperator agent, but are still valid cooperative strategies.
-Our results demonstrate that the use of a regret-based objective is critical to the success of our method.
+(2024) proposed training a Variational Autoencoder (Kingma & Welling, 2022), on simulated cooperative agent trajectories to generate synthetic training data to improve human-AI coordination.
+This approach outperforms previous methods because the generative model enables sampling random novel partner strategies that interpolate between or combine strategies from the data.
+GOAT leverages this approach to proactively generate challenging novel partners strategies to train the cooperator robustly.
 3 Preliminaries
-Two-player Markov game: We consider the two-player Markov game [19] as a tuple $\bigl{(}\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr{)}$ where $\mathcal{S}$ is the state space shared by all the agents.
-$\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ denote the action space of agents such that at each time step in the environment, each agent selects one action $a_{1,t}\in\mathcal{A}_{1}$, $a_{2,t}\in\mathcal{A}_{2}$ sampled from their own action space according to their policies $\pi_{1}:\mathcal{S}\to\Delta(\mathcal{A}_{1})$ and $\pi_{2}:\mathcal{S}\to\Delta(\mathcal{A}_{2})$.
+Two-player Markov game: We consider the two-player Markov game (Littman, 1994) as a tuple $\bigl(\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr)$ where $\mathcal{S}$ is the state space shared by all the agents.
+$\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ denote the action space of agents such that at each time step in the environment, and each agent selects one action $a_{1,t}\in\mathcal{A}_{1}$, $a_{2,t}\in\mathcal{A}_{2}$ sampled from their own action space according to their policies $\pi_{1}:\mathcal{S}\to\Delta(\mathcal{A}_{1})$ and $\pi_{2}:\mathcal{S}\to\Delta(\mathcal{A}_{2})$.
 After taking these actions, they receive a shared immediate reward $r:\mathcal{S}\times\mathcal{A}_{1}\times\mathcal{A}_{2}\to\mathbb{R}$.
@@
 The parameter $\gamma\in[0,1)$ is the discount factor that weights immediate versus future rewards.
-Given two policies $\pi_{1}$ and $\pi_{2}$, mapping each state to a distribution over possible actions for Agents 1 and 2, respectively, the expected return is captured by the value function $V(\pi_{1},\pi_{2})\;=\;\mathbb{E}\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr{]}$ which sums the discounted rewards from each timestep.
-We will use $V(\pi_{1},\pi_{2})$ in later sections to discuss how agents can maximize their joint return.
+The expected joint return is captured by the value function $V(\pi_{1},\pi_{2})\;=\;\mathbb{E}\Bigl[\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr]$ which sums the discounted rewards from each timestep.
 Generative model of cooperation strategies: Following Liang et al.
-[18], we train a Variational Auto-Encoder (VAE) [16] on diverse cooperative strategies.
-The VAE is trained to reconstruct coordination trajectories from a dataset $D$, which is obtained by using simulated PBT agents to interact with each other in the environment.
-Here, we use Maximum Entropy Population [43] to provide the simulated partner agents, since MEP was shown to provide the best performance in Liang et al.
-[18].
-After producing the dataset $D$ using MEP agents, we then train the generative model, specifically the VAE, on this data.
+(2024), we train a Variational Auto-Encoder (VAE) (Kingma & Welling, 2022) on diverse cooperative strategies.
+The VAE is trained to reconstruct the partner’s actions from a dataset of coordination trajectories $D$, which is obtained from interactions between simulated PBT (Population-Based Training) agents within the environment.
+Here, we use a fixed agent population with training methods like Maximum Entropy Population (Zhao et al., 2022) and CoMeDi (Sarkar et al., 2023) to provide the simulated partner agents.
 The VAE consists of an encoder $q(z|\tau;\phi)$, which maps a trajectory $\tau=(s_{0},a_{0},...,s_{T})$ to a latent space, where $\phi$ represents the parameters of the decoder.
 The decoder $p(a_{t}|z,\tau_{0..t-1};\phi)$ predicts actions $a_{t}$ autoregressively using the latent $z$ and the agent’s history $\tau_{0..t-1}$.
-The VAE model is optimized using the ELBO loss described in [16] as
-$\mathcal{L}(\theta,\phi)=\mathbb{E}_{\tau\sim D}\Bigg{[}\mathbb{E}_{z\sim q(\cdot|\tau,\phi)}\left[\sum_{t=1}^{T}\log p(a_{t}|z,\tau_{0..t-1};\phi)\right]-\beta\cdot KL\big{(}q(z|\tau,\phi)\,\|\,\mathcal{N}(0,I)\big{)}\Bigg{]}$
-The $\beta$ parameter balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space resembles a standard normal distribution $\mathcal{N}(0,I)$.
-This is important, because we can then sample new cooperation partners by sampling a $z\sim\mathcal{N}(0,I)$, and using the decoder to obtain a partner policy $\pi_{P}:=p(a_{t}|z,\tau_{0..t-1};\phi)$.
-The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations [1].
-During training, the $\beta$ term was scheduled to automatically save different versions of KL-regularized VAE models.
-This method by Liang et al.
-[18] has proved efficient in capturing diverse strategies and enabling strategy generation that is valid and cooperative.
+The VAE model is optimized using the ELBO loss described in (Kingma & Welling, 2022).
+The $\beta$ parameter that balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space resembles a standard normal distribution $\mathcal{N}(0,I)$.
+This is important for our method, because it enables generating new cooperation partners by sampling a latent $z\sim\mathcal{N}(0,I)$, and using the decoder to obtain a partner policy $\pi_{P}:=p(a_{t}|z,\tau_{0..t-1};\phi)$.
+The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations (Burgess et al., 2018).
 4 GOAT: Generative Online Adversarial Training
 We introduce GOAT, which combines generative modeling with online regret-based adversarial training to enable robust human-AI coordination.
-Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods [4].
-Here, an Adversary policy $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it.
+Adversarially minimizing Cross-Play (XP) performance.
+We begin by introducing an adversarial training objective based on the XP performance of the Cooperator $\pi_{C}$ and an adversarial partner $\pi_{A}$.
+The adversary $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it.
 This type of optimization is widely known as minimax optimization in game theory:
@@
 The Adversary can thus search over a well-structured, continuous latent space, allowing a smooth optimization landscape.
-We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation 1.
+We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation in Eq.
+1.
 Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce an Adversary embedding $z^{\prime}=\pi_{A}(z)$.
 This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$.
-Thus, the optimization problem is as follows:
+Let $\pi_{P}^{\pi_{A}(z)}$ be the partner policy produced by having the adversary $\pi_{A}$ choose an embedding which is then simulated with the generative model.
+Then, the optimization problem is as follows:
 $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}\left[\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},\pi_{C}(a^{C}_{t}|\tau_{0:t-1}),p(a^{P}_{t}|\pi_{A}(z),\tau_{0:t-1};\theta))\right]$ (2)
-$=$ $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}[V(\pi_{C},\pi_{P})]$ (3)
-However, this minimax objective still does not guarantee effective exploration in the partner space as it optimizes for worst-case returns.
+$=$ $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}[V(\pi_{C},\pi_{P}^{\pi_{A}(z)})]$ (3)
+This objective no longer directly trains the partner policy on the adversarial objective, but instead uses the adversary to steer the pre-trained generative model to select partners.
+Regret-based adversarial cooperative training.
+Because Eq.
+3 is still fundamentally a minimax objective, it does not guarantee effective exploration in the partner space because it optimizes for worst-case returns.
 Therefore, the Adversary will only search for the worst agents that the generative model can simulate.
-Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies which could have good performance, but the Cooperator failed to cooperate well with them.
-Regret-based adversarial cooperative training.
-To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent.
-We now formalize the definition of Regret we will use for cooperative settings.
-Self-Play (SP): This play style, first used by Samuel [27] to play checkers, is defined as the expected return when an agent is paired with itself:
-$J_{SP}=\mathbb{E}[V(\pi_{P},\pi_{P})]$ (4)
-Regret ($R$): As proposed initially by Robbins [25], regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight.
-We formalize regret in the cooperative setting as the performance gap between the cross-play (XP) performance an agent achieves with a partner $\pi_{P}$, and $\pi_{P}$’s optimal self-play (SP) performance.
+Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies that could have good performance, but for which the Cooperator still fails to perform well.
+Thus, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent.
+As proposed initially by Robbins (1952), regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight.
+We formalize regret in the cooperative setting as the performance gap between the cross-play (XP) performance an agent achieves with a partner $\pi_{P}$, and $\pi_{P}$’s optimal self-play (SP) performance, $\mathbb{E}[V(\pi_{P},\pi_{P})]$, which is the score it achieves when paired with itself.
 $Regret(\pi_{P},\pi_{C})=\mathbb{E}[V(\pi_{P},\pi_{P})-V(\pi_{P},\pi_{C})]=J_{SP}-J_{XP}$ (4)
-At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$.
+GOAT objective.
+At the start of the training process, GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$.
 This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$.
@@
 The full objective of the GOAT game is thus:
-$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
-The online optimization as shown in algorithm 1 ensures that the Adversary is incentivized to continually discover new partner strategies that are challenging for the Cooperator, by searching the latent space to maximize regret.
-The Cooperator in turn learns to adapt to these partner strategies by minimizing the regret.
+$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P}^{\pi_{A}(z)},\pi_{C})$ (5)
+The online optimization, as shown in algorithm 1, ensures that the Adversary is incentivized to continually discover new partner strategies that are challenging for the Cooperator, by searching the latent space to maximize regret.
+The Cooperator, in turn, learns to adapt to these partner strategies by minimizing the regret.
 This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones.
@@
 5 Experiments
-In this section, we describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
-We evaluate GOAT using two popular coordination benchmarks: 1)1)1 ) One-Step Cooperative Matrix Game (CMG) [20, 43, 4] and 2)2)2 ) Overcooked [3].
-Figure 2: Carroll et al.
-[3] first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit.
-To increase coordination complexity, Liang et al.
-[18] introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup.
-This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
-5.1 One-Step Cooperative Matrix Game (CMG):
-This is a single-step game where two agents (Row agent and Column agent) independently select an action by choosing a row and a column, respectively, and the intersection of their choices returns the associated reward according to the payoff matrix.
-Both agents receive the same reward and attempt to maximize the common payoff.
-A typical payoff matrix in CMG is defined by $(M,\{k_{m}\},\{r_{m}\})$ where $M$ is number of solutions, $k_{m}$ is number of compatible actions, and $r_{m}$ is the reward associated with each solution.
-We use a challenging multimodal layout called CMG-M as shown in Figure 3.
-5.2 Overcooked:
-This is a collaborative game where two players need to collaborate by dividing cooking tasks and avoiding obstructing each other.
+Our experimental evaluation is designed to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
+Environments.
+We evaluate GOAT using three popular coordination benchmarks: $1)$ One-Step Cooperative Matrix Game (CMG) (Lupu et al., 2021; Zhao et al., 2022; Charakorn et al., 2023), $2)$ Cooperative Reaching Game (CRG) (Rahman et al., 2023; 2024; Charakorn et al., 2023), and $3)$ Overcooked (Carroll et al., 2020).
+(a) CMG Payoff Matrix
+(b) Self-Play
+(c) MEP
+(d) CoMeDi
+(e) GAMMA
+(f) MiniMax
+(g) GOAT
+(h) Expected Reward
+Figure 2: $a)$ Cooperative Matrix Game.
+$b)\text{ to }g)$ Policy probability distribution to show coverage of different methods on the CMG payoff matrix $h)$ Total expected rewards for each method, assuming we uniformly sample partners and each method gets a payoff for each partner proportional to the amount of coverage they have for the reward block.
+  • One-Step Cooperative Matrix Game (CMG).
+CMG is a single-step game where two agents (Row and Column) independently select an action by choosing a row and a column, respectively, and the intersection of their choices returns the associated reward according to the payoff matrix.
+Both agents receive the same reward.
+A typical payoff matrix in CMG is defined by $(M,\{k_{m}\},\{r_{m}\})$ where $M$ is the number of solutions, $k_{m}$ is the number of compatible actions, and $r_{m}$ is the reward associated with each solution.
+We use a challenging layout shown in figure 2(a) with three scenarios: no reward, suboptimal rewards, and global maximum reward.
+  • Cooperative Reaching Game (CRG).
+CRG is a $5\text{x}5$ grid world, with $|\mathcal{A}|=5$, where two agents need to reach the same goal location simultaneously to receive rewards.
+For agents to cooperate and receive rewards, they must reach and stay in any of the same goal coordinates.
+An ideal Cooperator should chase their partner to the same goal or persuade them to follow the same goal.
+We evaluate GOAT and prior baselines against $11$ heuristic agent teammates described in detail in appendix E.
+  • Overcooked.
+A state-of-the-art collaborative, real-time benchmark where two players need to cooperate by dividing cooking tasks and avoiding obstructing each other.
 Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively.
-We test GOAT on the first five layouts introduced in [3], and add a more complex layout introduced by Liang et al.
-[18], the Multi-Strategy Counter (all shown in Figure 2).
-5.3 GOAT implementation & Baselines:
-The Cooperator is trained using PPO, and the Adversary is trained using REINFORCE.
-Training the Adversary is RL-algorithm-agnostic as it does not depend on state and is a one-step optimization process.
-Further details on the KL-regularization of the VAE, training procedure for GOAT, and other architecture and hyperparameters are available in the Appendix.
-We compare to 5 competitive prior baselines on human-AI coordination: Behavior Cloning (BC) partners with RL cooperator [3] (BC+RL), PBT methods Fictitious Co-Play (FCP) [32], maximum entropy population (MEP) [43], CoMeDi [28], and GAMMA trained on MEP (MEP+GAMMA) [18], which previously showed state-of-the-art performance in real human evaluations.
-5.4 Human Evaluation:
-We conducted a user study with 40 participants to discover which method best coordinated with humans and is evaluated as qualitatively better.
-Because we want the most challenging tasks with the most diverse strategy space, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter.
-The latter presents a more complex cooperation task because partners can prepare two dishes using different ingredients and thus represents a more rigorous test of whether algorithms can generalize to cooperate with real humans.
+We test GOAT on the most challenging tasks with a diverse strategy space.
+Counter Circuit introduced in Carroll et al.
+(2020), and a more complex layout introduced by Liang et al.
+(2024), Multi-Strategy Counter.
+GOAT implementation & Baselines: Training the Adversary is RL-algorithm-agnostic as it does not depend on state and is a one-step optimization process.
+To keep the optimization loop simpler, the Cooperator is trained using PPO, and the Adversary is trained using REINFORCE.
+Details on the KL-regularization of the VAE, training procedure for GOAT, and other architecture and hyperparameters are available in the Appendix.
+We compare to 5 competitive prior baselines on human-AI coordination: Behavior Cloning (BC) partners with RL cooperator (Carroll et al., 2020) (BC+RL), PBT methods Fictitious Co-Play (FCP) (Strouse et al., 2022), maximum entropy population (MEP) (Zhao et al., 2022), CoMeDi Sarkar et al.
+(2023), and GAMMA trained on MEP (MEP+GAMMA) Liang et al.
+(2024), which previously showed state-of-the-art performance in real human evaluations.
+Human Evaluation: We conducted a user study with 40 participants to discover which method best coordinated with humans in a real-time evaluation in Overcooked, recording the team scores.
 Participants were recruited from the Prolific online crowdsourcing platform following an IRB-approved protocol.
@@
 6 Experimental Results
-(a) CMG Layout
-(b) Self-play
-(c) TrajeDi
-(d) MEP
-(e) CoMeDi
-(f) GOAT
-Figure 3: To test the ability of the method to discover and exploit multiple minima and maxima in the problem space, we introduce a custom CMG layout (a).
-This reward structure is designed to have three scenarios: no reward, suboptimal rewards, and global maximum reward.
-The policies for the methods were aggregated and normalized into a heatmap.
-6.1 RQ1: Can GOAT discover and exploit multiple cooperative modes?
-For GOAT to be effective, the VAE should be able to discover and exploit regions of the reward space.
-To test this, we use a one-step cooperative matrix game (CMG).
-We test the cooperative policy generation capability of VAE by training it on trajectory data generated by the population of CoMeDi agents.
-Since CoMeDi data is cooperative, we hypothesize that the VAE should only generate policies that are in distribution.
-To keep the evaluation fair, we used a population size of 8 for all the methods tested on CMG-M.
-During training, GOAT covers all the modes and correctly assigns importance to the agent’s actions.
-As seen in Figure 3, the baselines self-play, Trajedi, and MEP get stuck and might not discover other modes.
-CoMeDi discovers all the regions and gives the highest priority to the global maximum, but spreads probability mass roughly equally across all other modes, regardless of their payoff.
-In contrast, GOAT not only covers all modes but is better at assigning higher probability to more rewarding modes (for example, assigning low probability to modes 4 and 7, which have low payoff).
-We note that the quality of VAE training and the amount of training data greatly influence the resulting GOAT policy.
-To robustly train the VAE, we follow Liang et al.
-[18] and augment the 8 policy pool data by pairing the row and column agents of one policy with other policies.
-In our experiments, we saw that higher $\beta$ for KL regularization in VAE resulted in better generation of actions.
-6.2 RQ2: Will adversarial training result in a more robust Cooperator with better sample complexity?
-PBT methods, which rely on simulated populations of self-play partners, are often computationally expensive and have poor coverage of actual human behaviors.
-Training a Cooperator against them may not achieve good performance with people, and it may take a long time since the training does not focus on correcting weaknesses in the Cooperator but instead randomly samples from the population.
+Figure 3: CRG: average reward obtained against the 11 Heuristic Agent teammates across 5 seeds of each method.
+RQ1: Can GOAT discover and exploit multiple cooperative strategies?
+For GOAT to be effective at generalizing to novel partners, the adversary should be able to discover and exploit all rewarding cooperative modes.
+We first test this question with the CMG environment, using a population size of 8 for all baseline PBT methods.
+To train the VAE using the population of CoMeDi agents, we use the same procedure as GAMMA Liang et al.
+(2024), which pairs each of the 8 row and column agents in the population together to generate trajectory data.
+Since VAE does not have a population, it can generate many more agents than PBT methods.
+As seen in Figure 2(b) and 2(c), the baselines’ Self-Play and MEP fail to cover all modes, implying they would not generalize to partners that used those modes.
+CoMedi 2(d) and GAMMA 2(e) discover all the regions and give the highest priority to the global maximum, but spread probability mass roughly equally across all other modes, regardless of their payoff.
+Whereas MiniMax 2(f) converges to the worst-case scenario.
+In contrast, GOAT not only covers all the modes but also is better at assigning higher probability to more rewarding modes (for example, assigning low probability to modes 3, 4, and 7, which have low payoff).
+Thus, GOAT achieves good coverage while also focusing on maximizing return.
+This is also evident in Figure 3, which shows the results of evaluating each method in the CRG environment.
+Here we see that GOAT achieves the highest average score when teamed with all 11 heuristic partner agents.
+(a) Counter Circuit
+(b) Simulation Results
+(c) Human evaluation
+(d) Multi-Strategy Counter
+(e) Simulation Results
+(f) Human evaluation
+Figure 4: Overcooked: Carroll et al.
+(2020) first introduced the most challenging Counter Circuit layout $(a)$.
+To increase coordination complexity, Liang et al.
+(2024) introduced the Multi-Strategy Counter layout $(d)$, where players could choose between tomato and onion ingredients to cook soup.
+This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
+$(b)\text{ \& }(e)$ are evaluation of GOAT against a human proxy model trained with BC.
+We compare to 4 baselines, which are trained using simulated population data, including the previous state-of-the-art method, GAMMA (Liang et al., 2024).
+Error bars show the std.
+err.
+over 5 random seeds.
+$(c)\text{ \& }(f)$ shows the evaluation of the performance of methods when tested against real humans in two layouts: counter circuit and multi-strategy counter, respectively.
+RQ2: Will adversarial training result in a more robust Cooperator with better sample complexity in simulation?
+PBT methods rely on simulated populations of self-play partners, are often computationally expensive, and have poor coverage of actual human behaviors.
+Training a Cooperator against them may not achieve satisfactory performance with people, and it may take a long time to converge since the training does not focus on correcting weaknesses in the Cooperator, but instead randomly samples from the population.
 We hypothesize that our adversarial training technique can more efficiently search the space of partners to make an agent that is more robust and requires less training time.
-To validate this hypothesis, we investigate the learning curves of all simulated training methods, including our own, in Figure 4.
-Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work [3, 32, 18].
+To validate this hypothesis, we investigate the Overcooked learning curves of all simulated training methods, including our own, in Figure 4(b) and 4(e).
+Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work (Carroll et al., 2020; Strouse et al., 2022; Liang et al., 2024).
 The human proxy enables the automatic evaluation of agents’ performance throughout training.
-As is evident in the curves, GOAT achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques.
-We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations.
-Adversarial exploration of this space enables rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies.
-By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to diverse partners.
-Because it efficiently targets training to weak points in the Cooperator’s policy, it has lower sample complexity than traditional PBT methods.
-Figure 4: Evaluation of GOAT against a simulated human policy trained with behavior cloning (BC).
-We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method, GAMMA [18].
-Error bars show the standard error over 5 random seeds.
-6.3 RQ3: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans?
-As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-strategy Counter layouts.
-We recruited novel human players whom the agents had not previously seen.
-Figure 5 depicts the evaluation results for the two layouts and all the agents against human players.
-On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA (see Table 4 for statistical significance results).
-We hypothesize this is because both methods are reaching close to optimal performance on this layout.
-However, on the more complex layout, Multi-Strategy Counter, the GOAT leads to significantly higher cooperation performance with real humans, leading to a 38% improvement over the previous state of the art.
-To test these differences for yourself, please visit our webpage (https://sites.google.com/view/goat-2025/), which provides an interactive demo in which you can play Overcooked with both GOAT and the baselines.
-Figure 5: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit (left plot) and multi-strategy counter (right plot).
-6.4 RQ4: Will adversarial training create a curriculum of increasingly challenging strategies by discovering them in latent space?
+As is evident in the curves, GOAT achieves significantly higher performance than all 4 baselines in both Overcooked layouts and increases performance more quickly than other techniques.
+We hypothesize this is because the generative model, trained on cooperative trajectories, encodes a latent space rich in strategic variations.
+Adversarial exploration in this latent space enables rapid coverage of diverse human behaviors ranging from simple coordination strategies to complex, context-dependent strategies.
+By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies and exposes the Cooperator to all the semantically diverse strategies that are beneficial for generalizing to diverse partners.
+It efficiently targets training to weak points in the Cooperator’s policy, and thus has lower sample complexity than traditional PBT methods.
+RQ3: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans?
+The real test of a human-AI coordination algorithm is whether it can work with actual people.
+As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-Strategy Counter Overcooked layouts.
+We recruited novel human players with no prior exposure to the agents.
+Figure 4(c) and 4(f) depict the evaluation results for the two layouts and all the agents against human players.
+GOAT outperforms the previous baselines on both layouts.
+Even though GOAT shows significant improvement over baselines on Counter Circuit simulation results, the magnitude of the improvement over GAMMA is smaller ($3\%$) on this layout.
+This is because both methods are reaching close to optimal performance in this simpler environment.
+However, on the more complex layout, Multi-Strategy Counter, GOAT leads to markedly higher cooperation performance with real humans, with a 38% improvement over the previous state of the art.
+We encourage readers to test these differences for themselves by visiting our webpage, which provides an interactive demo in which you can play Overcooked with both GOAT and the baselines https://sites.google.com/view/goat-2025/home.
+(a) GOAT exploring the latent space
+(b) Regret vs Minimax Exploration
+Figure 5: $a)$ GOAT is exploring the latent space over training episodes, compared to red standard normal distribution cluster of GAMMA.
+$b)$ Comparison of Regret (Blue) with Minimax (Red) in the latent space.
+RQ4: Will adversarial training create a curriculum of increasingly challenging strategies by discovering them in latent space?
 The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges.
@@
 This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods.
-Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
-Figure 6: Left: GOAT exploring the latent space exploration compared against GAMMA [18] in red over the whole training period.
-Right: Minimax Adversary in red and GOAT in blue exploring the latent space over the training period.
-We provide a visual example related to this on our webpage.
-The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 6.
-We compare this to the distribution of embeddings of the original GAMMA method [18] samples throughout training (red ball).
-We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple).
+5(a) assesses how GOAT explores the latent space of the generative model over the whole training interval.
+Each point is a latent vector, and the colorbar represents the training iteration number.
+The 5(a) compares GOAT embeddings to the distribution of embeddings of the original GAMMA method (Liang et al., 2024) samples throughout training (red ball).
+We can see that at the start of training, GOAT samples latent vectors distributed throughout this ball (shown in purple).
 As training progresses, it chooses a particular region to exploit.
@@
 Finally, at the end of the training step in yellow, GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
-6.5 RQ5: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective?
-In Section 4, we described how the purely adversarial minimax objective (Equation 1) is misaligned with effectively training agents to cooperate.
-In Equation 3, we showed how integrating a generative model could help alleviate this problem.
-However, while the generative model can prevent sabotage behavior, that does not mean all agents it simulates will show productive cooperation behavior.
-We claimed that the purely adversarial minimax objective would still focus too much on poorly performing or incompetent partners, motivating the use of the regret objective in Equation 5.
-We now empirically support this claim by comparing GOAT to an ablation of our method, which uses the minimax objective in Equation 3, on the Multi-Strategy Counter layout.
-As we can see in Figure 7, the GOAT outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT.
-We further investigate the embeddings chosen by both types of Adversary throughout the training period in Figure 6 (right).
-We find that the minimax Adversary mostly fixates on a single region (red cluster in the plot) where the policies result in low reward behaviors.
-One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance, but makes no attempts to cooperate.
-In the blue, we have embeddings chosen by the GOAT throughout the training period.
-We can observe that the regret objective causes the GOAT to move to explore multiple modes over the course of training, exploring the latent space more effectively and thus covering a wider range of interesting behaviors.
+RQ5: Ablation study of Minimax vs.
+Regret.
+In Section 4, we hypothesized that even with a generative model in the loop, the purely adversarial minimax objective (Eq.
+3) is misaligned with effectively training agents to cooperate.
+While the generative model can prevent sabotage behavior, because minimax focuses on worst-case scenarios, we hypothesized that a minimax adversary would still focus too much on poorly performing or incompetent partners.
+In this experiment, we directly compare training the GOAT adversary with either minimax (Eq.
+3) or regret (Eq.
+4) on the Multi-Strategy Counter Overcooked layout.
+We further investigate the embeddings chosen by both types of adversary throughout the training period in 5(b).
+We find that the minimax adversary mostly fixates on a single region (red cluster in the plot) where the policies result in low-reward behaviors or unproductive behaviors.
+One such example, shown as a video on our website, is where the minimax agent avoids working with its partner and converges to doing just enough not to sabotage the task.
+The team still receives rewards as the partner policy keeps completing the task uninterrupted.
+In contrast, we see that the regret objective causes GOAT to move to multiple modes in the embedding space over the course of training, exploring the latent space more effectively and thus covering a wider range of interesting behaviors.
+As shown in game videos on our website, it starts by doing one task, then shifts to another task, and then does both tasks simultaneously.
+We can also see robust training instances like intentionally coming in the way or taking roles.
 7 Conclusion
@@
 Thus, we enable robust learning without the risk of degenerate strategies.
-Experimental results on the Overcooked benchmark show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners.
-The improved coverage of the partner space demonstrated by our agents suggests this approach successfully bridges the gap between training efficiency and the generalization to human behavior.
+Experimental results on three cooperation tasks show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners.
+The improved coverage of the partner space demonstrated by our agents suggests this approach can successfully and robustly achieve generalization to human behavior.
 Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training.
-8 Acknowledgements
-This work was generously supported by the Cooperative AI Foundation, UW-Amazon Science Gift Hub, Amazon Middle Mile Products and Technology Team, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), and Microsoft Accelerate Foundation Models Research Program.
+8 Acknowledgments
+This research was supported by the Cooperative AI Foundation, the UW-Amazon Science Gift Hub, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), the Microsoft Accelerate Foundation Models Research Program, Character.AI, DoorDash, and the Schmidt AI2050 Fellows program.
+This material is based upon work supported by the Defense Advanced Research Projects Agency and the Air Force Research Laboratory, contract number(s): FA8650-23-C-7316.
+Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of AFRL or DARPA.
 Appendix A Reproducibility
-We have our live interactive demo and extended experimental data on this webpage.
-https://sites.google.com/view/goat-2025.
-We provide all the details to set up the environment and reproduce the results.
-For the One-Step Cooperative Matrix Game please find Google Colab with code and all the instructions to run the code.
+We have our live interactive demo, extended experimental data, and code on our webpage, link on first page.
 Appendix B Implementation details
@@
 Due to a smaller amount of data, we used data augmentation to train the VAE.
-Cooperator Agent The Cooperator agent was trained using PPO [29].
+Cooperator Agent The Cooperator agent was trained using PPO (Schulman et al., 2017).
 To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up.
-We adapted our codebase on the previous implementation of HSP [41] and GAMMA [18].
-GOAT For the GOAT policy training, we use the REINFORCE objective with KL divergence penalty for drifting away from the prior, mean, and standard deviation hard bounds to keep the policy constrained in the VAE latent space.
+We adapted our codebase on the previous implementation of HSP (Yu et al., 2022) and GAMMA (Liang et al., 2024).
+GOAT For GOAT policy training, we use the REINFORCE objective with KL divergence penalty to keep the policy constrained in the VAE latent space.
 This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
 Appendix C Hyperparameters
-The GOAT uses a simple policy and was trained using REINFORCE to generate latent vectors. The architecture and hyperparameters are defined below.
+GOAT uses a simple policy and was trained using REINFORCE to generate latent vectors. The architecture and hyperparameters are defined below.
 hyperparameter value
@@
 activation function ReLU
-output dimension 2×2$\times$2 × z_dim
+output dimension $2\times$ z_dim
 optimizer Adam
-learning rate 0.001
+learning rate 0.0005
 weight decay 0.0001
 latent dimension (z_dim) 16
-KL coefficient 0.1
-entropy coefficient 0.01
-std scaling 4.0, 1.0
-mean scaling 60.0
-gradient clipping 1.0
-learning rate scheduler start factor 1.0
-learning rate scheduler end factor 0.9
+KL coefficient 5.0
 Table 1: Hyperparameters for GOAT Generator
@@
 One Cooperator agent can be trained in roughly a day.
-The primary tests require roughly 3600 GPU hours.
+The primary tests require roughly 24-36 GPU hours for Overcooked and 1 hour for rest of the environments.
 We perform some preliminary experiments to determine the best training frameworks and hyperparameters.
-Appendix E Minimax and Regret comparison on overcooked
-Figure 7: Performance comparison between minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner.
-Error bars are the Standard Error of the Mean.
-Appendix F Cooperative strategies learned by VAE
-Figure 8: Left: We collect 4 cooperative policies on CMG-M using CoMeDi and then aggregate them.
-Right: The policies are used for training the VAE, where the latent dimension is 2 and we estimate the policy reconstruction from the VAE.
-Figure 9: After training a VAE on a 4-population pool of CoMeDi agents, we visualize the generation by interpolating the dimensions of the latent vector.
-Particularly for this example, we choose the latent dimension to be 2.
-We can see that VAE trained on cooperative policies has only encoded those policies in the latent space.
-As we move from one side of the plot to the other, we can see that different cells are activated, and it covers the cooperative policies as seen in the Figure 8
-Appendix G Bonferroni Tests
-Hypothesis Counter circuit ($p$-value) Multi-Strategy Counter ($p$-value)
-CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
-FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
-BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
-MEP $1.0$ $3.48\times 10^{-28}$
-MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
-[…]
-Appendix H Tests on CMG-S
+Appendix E Cooperative Reaching Game: Game layout & 11 Heuristic Agents
+Figure 6: Cooperative Reaching Game (CRG): Red and Blue are the two agents that should cooperate to reach and stay on one of the goal coordinates in the corner.
+Yellow goal coordinates lead to a reward of 11, and Green goal coordinates lead to a reward of $0.75$.
+  • Heuristic H01: Selects actions moving teammates toward the nearest reward-providing coordinate.
+  • Heuristic H02: Selects actions moving teammates toward the reward-providing coordinate most distant from their episode starting position.
+  • Heuristic H03: Moves teammates toward the nearest optimal reward-providing coordinate.
+  • Heuristic H04: Moves agents toward the optimal reward-providing coordinate most distant from teammates’ initial episode location.
+  • Heuristic H05: Same as H4, but considers only suboptimal reward-providing coordinates instead of optimal ones.
+  • Heuristic H06: Same as H5, but teammates move toward the nearest suboptimal reward-providing coordinate.
+  • Heuristic H07: At episode start, agents randomly select a reward-providing coordinate and move toward it.
+  • Heuristic H08: Moves teammates toward the reward-providing coordinate nearest to their counterpart agent’s location.
+  • Heuristic H09: Same as H8, but considers only optimal reward-providing coordinates as destinations.
+  • Heuristic H10: Moves teammates directly toward their counterpart agent’s location.
+  • Heuristic H11: Always randomly selects actions from teammates’ available action set.
+Appendix F Tests on CMG-S
 (a) CMG-S layout
@@
 GOAT successfully assigns priorities to regions.
+Appendix G Minimax and Regret comparison on overcooked
+Figure 8: Performance comparison between minimax vs GOAT objective on Multi-Strategy Counter when evaluated against a held-out human behavior cloned model as a partner.
+Error bars are the Standard Error of the Mean.
+GOAT outperforms MiniMax, but here MiniMax produces poor-performing agents that fail to participate in coordination.
+See game videos on our website.
+Appendix H Cooperative strategies learned by VAE
+Figure 9: Left: We collect 4 cooperative policies on CMG-M using CoMeDi and then aggregate them.
+Right: The policies are used for training the VAE, where the latent dimension is 2 and we estimate the policy reconstruction from the VAE.
+Figure 10: After training a VAE on a 4-population pool of CoMeDi agents, we visualize the generation by interpolating the dimensions of the latent vector.
+Particularly for this example, we choose the latent dimension to be 2.
+We can see that VAE trained on cooperative policies has only encoded those policies in the latent space.
+As we move from one side of the plot to the other, we can see that different cells are activated, and it covers the cooperative policies as seen in the Figure 9
 Appendix I Broader Impact Statement
@@
 GOAT has the potential to improve capabilities for household robots, vehicles, and collaborative AI systems where seamless human interaction adds value to people’s lives.
-This type of technology enhances human modelling, which can be deployed widely to help people from all walks of life and of all age groups.
+This type of technology enhances human modeling, which can be deployed widely to help people from all walks of life and of all age groups.
 We would also like to note that the development of such an AI system will accelerate technological progress and requires responsible development and deployment.
-Appendix J Human qualitative ratings.
-(a) Counter circuit layout
-(b) Multi-strategy Counter
-Figure 11: Human evaluations of agent behavior.
-Participants rated their agreement with the statement "The agent adapted to me when making decisions" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree.
-GOAT shows decent adaptability on simpler layout, whereas it appears more adaptable on harder layout.
-(a) Counter circuit layout
-(b) Multi-strategy Counter
-Figure 12: User assessment of agent coordination.
-Participants rated their agreement with the statement "The agent’s actions were human-like" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree.
-GOAT is seen as more confident in coordinating on harder layouts.
-(a) Counter circuit layout
-(b) Multi-strategy Counter
-Figure 13: User assessments of agent frustration levels.
-Participants rated their agreement with the statement "The agent’s behavior was frustrating" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree.
-GOAT receives comparable ratings for showing less frustrating behavior compared to their respective baselines.
-Appendix K Limitation
-The effectiveness of this method heavily depends on the training of VAE models and enough data available for VAE to learn the distribution instead of overfitting to a few examples.
+Appendix J Limitation
+The effectiveness of this method depends on the training of VAE models and enough data available for VAE to learn the distribution, instead of overfitting to a few examples.
 It would be a promising method for fields like robotics and human-AI interaction, where data availability is not an issue.